### v2

Abstract
Scene Graph Generation (SGG) aims to detect all the visual relation triplets $<$sub, pred, obj$>$ in a given image. With the emergence of various advanced techniques for better utilizing both the intrinsic and extrinsic information in each relation triplet, SGG has achieved great progress over the recent years. However, due to the ubiquitous long-tailed predicate distributions, today’s SGG models are still easily biased to the head predicates. Currently, the most prevalent debiasing solutions for SGG are re-balancing methods, e.g., changing the distributions of original training samples. In this paper, we argue that all existing re-balancing strategies fail to increase the diversity of the relation triplet features of each predicate, which is critical for robust SGG. To this end, we propose a novel Multi-level Compositional Feature Augmentation (MCFA) strategy, which aims to mitigate the bias issue from the perspective of increasing the diversity of triplet features. Specifically, we enhance relationship diversity on not only feature-level, i.e., replacing the intrinsic or extrinsic visual features of triplets with other correlated samples to create novel feature compositions for tail predicates, but also image-level, i.e., manipulating the image to generate brand new visual appearance for triplets.
Due to its model-agnostic nature, MCFA can be seamlessly incorporated into various SGG frameworks. Extensive ablations have shown that MCFA achieves a new state-of-the-art performance on the trade-off between different metrics.
Keywords:
Unbiased Scene Graph Generation Feature Augmentation Category ClusteringCompositional Learning
1 Introduction
As one of the fundamental comprehensive visual scene understanding tasks, Scene Graph Generation (SGG) has attracted unprecedented interest from our community and has made great progress in recent years Lu et al. (2016); Zhang et al. (2017); Zellers et al. (2018); Chen et al. (2019a, b); Tang et al. (2020); Li et al. (2021b, 2022a). Specifically, SGG aims to transform an image into a visually-grounded graph representation (i.e., scene graph) where each node represents an object instance with a bounding box and each directed edge represents the corresponding predicate between the two objects. Thus, each scene graph can also be formulated as a set of visual relation triplets (i.e., $<$sub, pred, obj$>$). Since such structural representations can provide strong explainable potentials, SGG has been widely-used in various downstream tasks, such as visual question answering, image retrieval, and captioning.
Figure 1: (a) The intrinsic and extrinsic information for SGG. The entity prediction is for the green box, and the predicate prediction is for the relation between the red and green boxes. (b) Illustration of the diversity of feature space and decision boundary between on and laying on before and after using re-balancing and MCFA. Each sample denotes the corresponding visual triplet features.
In general, due to the extremely diverse visual appearance of different visual relation triplets, recent SGG methods all consider both intrinsic and extrinsic information for entity and predicate classification Zellers et al. (2018); Tang et al. (2019); Lu et al. (2021).
By “intrinsic information”, we mean these intrinsic characteristics of the subjects and objects, such as their visual, semantic, and spatial features. For example in Fig. 1(a), with the help of these intrinsic features, we can easily infer all possible entity categories (e.g., pole or surfboard) and predicate categories (e.g., on or holding) of each triplet. However, sometimes it is still hard to confirm the exact correct predictions with only intrinsic information, especially for tiny objects. Thus, it is also essential to consider other “extrinsic information” in the same image, such as the context features from neighbor objects. As shown in Fig. 1(a), after encoding the features of surrounding objects (e.g., wave and hand), we can easily infer that the categories of the entity and predicate should be surfboard and holding.
Although numerous advanced techniques have been proposed to enhance SGG performance. Today’s SGG methods still fail to predict some informative predicates due to the ubiquitous long-tailed predicate distribution in prevalent SGG datasets Krishna et al. (2017). Such a distribution is characterized by a small amount categories with vast samples (head [^1] We directly use “tail”, “body”, and “head” categories to represent the predicate categories in the tail, body, and head parts of the number distributions of different predicates in SGG datasets, respectively.) and plenty of categories with rare samples (tail). As a result of the discrepancy of sample size among different categories, the learned decision boundary becomes improper (cf.. Fig. 1(b)), i.e., their predictions are biased towards the head predicates (e.g., on) and they are error-prone for the tail ones (e.g., laying on).
To overcome the bias issue, the most prevalent unbiased SGG solutions are re-balancing strategies, e.g., sample re-sampling Li et al. (2021b) and loss re-weighting Yan et al. (2020); Chen et al. (2022); Li et al. (2022c); Lyu et al. (2022). They alleviate the negative impact of long-tailed distribution by increasing samples or loss weights of tail classes. Then, the decision boundaries are adjusted to reduce the bias introduced by imbalanced distributions. However, we argue that all the existing re-balancing strategies only change the frequencies or contributions of existing relation triplets to improve the skewed decision boundaries, but fail to enhance the diversity of relation triplet features [^2] We use the “relation triplet feature” to represent the combination of both intrinsic and extrinsic features of each visual relation triplet. of each predicate, which is crucial to infer their complete feature space. This sparsity of feature diversity will lead to under-representation of tail categories, thereby making it hard to infer the complete data distribution, i.e., making it challenging to find the optimal direction to adjust the decision boundaries Chu et al. (2020); Vigneswaran et al. (2021). For example in Fig. 1(b), the feature space of laying on is so sparse that the decision boundary can be adjusted within a large range. The performance of this naive adjustment without “complete” distribution is always sensitive to hyperparameters, i.e., excessively increasing the sample number or loss weight of tail predicates may cause some head predicate samples to be incorrectly predicted as tail classes (right solid line), and vice versa (left solid line).
Figure 2: (a) Intrinsic-CFA: Replacing the entity feature of tail predicate triplet dog-laying on-bed from dog to cat to enhance the intrinsic feature. (b) Extrinsic-CFA: Mixing up the feature of tail predicate triplet pillow-laying on-bed into the context of pillow-on-bed to enhance the extrinsic feature.
Therefore, in this paper, we focus on enhancing diversity of relation triplet features instead of re-balancing their frequencies to learn more robust feature space and decision boundaries for predicates. One intuitive idea is to borrow novel entity features or context features from other existing triplet samples to replace the original features, thereby constructing new triplet features for the target predicates. By deeply exploring the correlations of objects and predicates, we can reasonably select the features to replace, ensuring the reliability of the re-composed relation triplet features. However, this naive augmentation technique on feature-level manipulation has its inevitable limitations. On the one hand, directly replacing the original visual features may lead to unnatural transitions (e.g., fusing a pillow’s feature with another one that are inadvertently entangled with adjacent objects, such as bedside cupboard) as shown in Fig. 3(b), further hurting the expressive ability of the model. On the other hand, the augmented samples are still limited to the in-domain distribution, and as a result, the trained models are unable to handle with out-of-domain predictions. Therefore, we also propose to manipulate relation triplet features on image-level with the help of powerful generative models as shown in Fig. 3(c).
Figure 3: (a) Context Image: The image which provides context information and is used to to augment with the triplet pillow-on-bed. (b) Feature-level Augmentation: Visualization of augmented results with the feature of tail predicate triplet pillow-laying on-bed selected from dataset. (c) Image-level Augmentation: Augmented new image with target triplet and context information by generative models.
In this paper, we propose a novel Multi-level Compositional Feature Augmentation (MCFA) strategy for unbiased SGG, which tries to solve the bias issue by enhancing the diversity of relation triplet features on both feature-level and image-level, especially for tail predicates.
As shown in Fig. 1(b), by enhancing the feature diversity of the tail predicates (e.g., laying on), SGG models can easily learn the proper decision boundaries (vs. re-balancing strategies).
On feature-level, we apply different strategies for augmenting intrinsic and extrinsic feature diversity respectively. For intrinsic diversity, we replace the entity features (e.g., subject or object) of a tail predicate triplet with other “suitable” entity features. To determine the suitable entity categories for the augmentation, we propose a novel relation-conditioned entity clustering method to find the predicate-aware correlations between different entity categories, and then we regard the entity features from the same cluster are suitable. Specifically, we calculate the category correlation by pattern, context, and semantic similarities. For example in Fig. 2 (a), the entity categories cat and dog are in the same cluster, i.e., we can augment the intrinsic feature by replacing the dog entity feature with a cat entity feature. For extrinsic diversity, we take advantage of the context or interactions of other triplets (i.e., context triplets) and enhance the features of tail predicate triplets by these context triplets. Specifically, given a context triplet randomly selected from an image, we first select a reasonable tail predicate triplet as the target by limiting the categories and relative position of two objects. Then, to minimize the impact on the prediction of other triplets in the original image and make use of the extrinsic features of the context triplet, we use an adaptive mixup strategy to fuse the features of targeted tail predicate triplet into the context triplet and adaptively adjust mixup weights based on predicate semantic similarity. For example in Fig. 2 (b), triplet pillow-laying on-bed is mixed up into the image of pillow-on-bed. To avoid introducing noise into augmented samples, we adaptively assign weights to the mixed triplets by considering the semantic similarity of relation triplets.
On image-level, we aim to directly manipulate the images and generate brand-new visual appearances for target predicate triplets. A core challenge in this process is ensuring the semantic plausibility and spatial coherence of the generated images. To address this, we first employ a relative position constraint to filter only contextually appropriate images associated with target triplet categories for augmentation. For the synthesis process, we leverage the powerful grounded text-to-image generation capabilities of recent generative models (e.g., GLIGEN Li et al. (2023b)), as shown in Fig. 3(c). Specifically, the model is conditioned on the original image, the target triplet’s categories, and corresponding entity grounding information to generate a new high-quality visual instance. Subsequently, we assess the quality of each generated image via a visual-language similarity score to ensure high fidelity and relevance. As the context triplets are selected from other predicate triplets and visual features of target triplets are novelly generated, both extrinsic and intrinsic diversity are enhanced under our image-level augmentation strategy.
We evaluate MCFA on two most prevalent and challenging SGG datasets: Visual Genome (VGKrishna et al. (2017) and GQA Hudson and Manning (2019). Since MCFA is a model-agnostic debiasing strategy, it can be seamlessly incorporated into various SGG architectures and consistently improve their performance. Unsurprisingly, MCFA can achieve a new state-of-the-art performance on the trade-off between different metrics. Extensive ablations and results on multiple SGG tasks and backbones have shown the generalization ability and effectiveness of MCFA.
This paper is a comprehensive and in-depth extension of our previous ICCV work Li et al. (2023a) with several significant improvements:
  1. 1.
    We propose a novel relation-conditioned entity clustering strategy for feature-level intrinsic augmentation by leveraging the predicate-driven semantic knowledge, leading to more accurate groupings.
  2. 2.
    We propose an adaptive mix-up strategy for feature-level extrinsic augmentation, which dynamically adjust mix-up coefficient based on predicate similarity, effectively handling with the varying semantic context of triplets.
  3. 3.
    We devise a novel generative augmentation method to manipulate relation triplet features on the image-level and generate high-quality samples, thereby effectively boosting the performance of tail predicates.
  4. 4.
    Extensive qualitative and quantitative analyses demonstrate the effectiveness of our MCFA.
2 Related Work
Unbiased Scene Graph Generation. Biased predictions prevent further use of scene graphs in real-world applications. Recent unbiased SGG works can be divided into three main categories: 1) Re-balancing: It alleviates the negative impact of long-tailed predicate distribution by re-weighting or re-sampling Yan et al. (2020); Li et al. (2022c); Lyu et al. (2022); Li et al. (2021b); Desai et al. (2021). 2) Unbiased Inference: It makes unbiased predictions based on biased models Tang et al. (2020); Yu et al. (2021). 3) Noisy Label Learning: It reformulates SGG as a noisy label learning problem and corrects the noisy samples Li et al. (2022a, b). This work points out the drawbacks of existing re-balancing methods and studies unbiased SGG from the new perspective of multi-level feature augmentation.
Feature Augmentation. Data augmentation is a prevalent training trick to improve models’ performance. Conventional data augmentation methods DeVries and Taylor (2017); Zhong et al. (2020) usually synthesize new samples in a hand-crafted manner. Compared to these image-level data augmentation methods, feature augmentation is another efficient way to improve models’ generalizability by directly synthesizing samples in the feature space Chu et al. (2020); Li et al. (2021a). Compared to existing data augmentation applications, SGG is a sophisticated task that involves intrinsic and extrinsic features. In this work, we propose MCFA to enrich the diversity of relation triplet features for debiasing from both feature-level and image-level.
Compositional Learning (CL). CL has been successfully applied to various computer vision tasks. As for visual scene understanding, some Human Object Interaction (HOI) detection works Hou et al. (2020); Kato et al. (2018); Hou et al. (2021) compose new interaction samples that significantly benefit both low-shot and zero-shot settings. Inspired by these successes, we adopt this compositional approach for unbiased SGG. Our method enhances triplet feature diversity by first decomposing representations into their intrinsic and extrinsic constituents, and then augmenting these two feature streams.
Figure 4: The illustration of unbiased SGG framework with MCFA on feature-level augmentation.
Figure 5: The pipeline of feature-level and image-level augmentation strategy. Feature-level strategy consists of two approaches to generate new compositions of triplet features tailored for intrinsic and extrinsic diversity enhancement respectively, i.e., (a) Intrinsic-CFA and (b) Extrinsic-CFA. While image-level strategy generates new visual appearance of relation triplet by generative models to ensure the augmentation quality, i.e.(c) Image-CFA.
3 Multi-level Compositional Feature Augmentation
Given an image $I$, a scene graph is formally represented as $\mathcal{G}$ = $\mathcal{N}$, $\mathcal{E}$}, where $\mathcal{N}$ and $\mathcal{E}$ denote the set of all objects and their pairwise visual relations, respectively. Specifically, the $i$-th object in $\mathcal{N}$ consists of a bounding box (bbox) $\bm{b}_{i}\in\mathcal{B}$ and its entity category $o_{i}\in\mathcal{O}$. A relation $r_{ij}\in\mathcal{R}$ denotes the predicate category between $i$-th object and $j$-th object. $\bm{b}_{ij}$ denotes the union box of bbox $\bm{b}_{i}$ and bbox $\bm{b}_{j}$. $\mathcal{B}$, $\mathcal{O}$, and $\mathcal{R}$ represent the set of all entity bboxes, entity categories, and predicate categories, respectively.
In this section, we first revisit the two-stage SGG baselines in Sec. 3.1. Then, we elaborate the pipeline of MCFA, including feature-level and image-level feature augmentation strategy (cf., Fig. 4) in Sec. 3.2 and Sec. 3.3, respectively. Finally, we detail the training objectives in Sec. 3.4.
3.1 Revisiting the Two-Stage SGG Baselines
Since the mainstream SGG methods are two-stage models, we review the two-stage SGG framework here Zellers et al. (2018); Tang et al. (2019). A typical two-stage SGG model involves three steps: proposal generation, entity classification, and predicate classification. Thus, the SGG task $P(\mathcal{G}|I)$ is decomposed into:
$P(\mathcal{G}|I)=P(\mathcal{B}|I)P(\mathcal{O}|\mathcal{B},I)P(R|\mathcal{O},\mathcal{B},I).$ (1)
Proposal Generation $P(\mathcal{B}|I)$. This step aims to generate all bbox proposals $\mathcal{B}$. Given an image $I$, they first utilize an off-the-shelf object detector (e.g., Faster R-CNN Ren et al. (2015)) to detect all the proposals $\mathcal{B}$ and their visual features $\{\bm{v}_{i}\}$.
Entity Classification $P(\mathcal{O}|\mathcal{B},I)$. This step mainly predicts the entity category of each $\bm{b}_{i}$in $\mathcal{B}$. Given a visual feature $\bm{v}_{i}$ and proposal $\bm{b}_{i}$, they use an object context encoder $\mathtt{Enc}_{obj}$ to extract the contextual entity representation $\bm{f}_{i}$:
$\bm{f}_{i}=\mathtt{Enc}_{obj}(\bm{v}_{i}\oplus\bm{b}_{i}),$ (2)
where $\oplus$ denotes concatenation. Then, they use an object classifier $\mathtt{Cls}_{obj}$ to predict their entity categories:
$\widehat{o}_{i}=\mathtt{Cls}_{obj}(\bm{f}_{i}).$ (3)
Predicate Classification $P(\mathcal{R}|\mathcal{O},\mathcal{B},I)$. This step predicts the predicate categories of every two proposals in $\mathcal{B}$ along with their entity categories. First, they use a relation context encoder $\mathtt{Enc}_{rel}$ to extract the refined entity feature $\bm{\tilde{f}}_{i}$:
$\bm{\tilde{f}}_{i}=\mathtt{Enc}_{rel}(\bm{v}_{i}\oplus\bm{f}_{i}\oplus\bm{w}_{i}),$ (4)
where $\bm{w}_{i}$ is the GloVe embedding Pennington et al. (2014) of predicted $\widehat{o}_{i}$. It is worth noting that both $\mathtt{Enc}_{cls}$ and $\mathtt{Enc}_{rel}$ often adopt a sequence model (e.g., Bi-LSTM Zellers et al. (2018), Tree-LSTM Tang et al. (2019), or Transformer Yu et al. (2021)) to better capture context. After relation feature encoding, they use a relation classifier $\mathtt{Cls}_{rel}$ to predict the relation $\widehat{r}_{ij}$ between any two subject-object pairs:
$\widehat{r}_{ij}=\mathtt{Cls}_{rel}([\bm{\tilde{f}}_{i}\oplus\bm{\tilde{f}}_{j}]\circ\bm{u}_{ij}),$ (5)
where $\circ$ denotes element-wise product, and $\bm{u}_{ij}$ denotes the visual feature of the union box $\bm{b}_{ij}$.
3.2 Feature-level CFA
In this paper, we treat a relation triplet feature2 in SGG as a combination of two components: intrinsic feature and extrinsic feature, which refer to the intrinsic information (i.e., the subject and object itself) and extrinsic information (i.e., the contextual objects and stuff), respectively. Correspondingly, our feature-level CFA consists of intrinsic-CFA and extrinsic-CFA to augment these two types of features. For ease of presentation, we call the “targeted tail predicate triplet whose feature to be enhanced” as “query triplet”, and the “relation triplet whose image is used to provide context” as “context triplet”. Besides, to facilitate the feature augmentation, we store all the visual features of all tail predicate triplets (i.e., visual features $\bm{v}$ of two entities and their union feature $\bm{u}$) in a feature bank before training. Besides, we adopt the repeat factor $\eta$ = $\max(1,\eta_{r})$ Gupta et al. (2019); Li et al. (2021b) to sample images to provide enough context triplet and query triplet for augmentation, where $\eta_{r}$ = $\sqrt{\lambda/{f_{r}}}$, $f_{r}$ is the frequency of predicate category $r$ on the entire dataset, and $\lambda$ is hyperparameter. The unbiased SGG pipeline with feature-level CFA is shown in Fig. 4.
3.2.1 Intrinsic-CFA
Intrinsic-CFA enhances the feature of query triplet by replacing its visual features of entities (i.e., subject and object). It consists of two steps: entity category selection and entity feature selection and replacement (cf.. Fig. 5(a)). During training, we randomly select a query triplet from a batch of images, and randomly select one of its entity features as input. Then we put it into the Intrinsic-CFA module (cf.. Fig. 4). Next, we detailedly introduce each step.
Entity Category Selection. This step is used for determining the category to which the entity feature of the query triplet is replaced. Firstly, we propose a novel relation-conditioned entity clustering that organizes entity categories based on their semantic role in a triplet. The framework consists of the following procedures.
(1) Predicate Clustering. To maintain consistency in entity substitutions, we first group predicates into several clusters. The clustering process utilizes three similarity measures:
  • Pattern Similarity: This metric quantifies the degree of functional overlap between predicates by analyzing the frequency with which they govern similar subject-object pairs. Formally, for two predicate categories $c_{p_{i}}$ and $c_{p_{j}}$, pattern similarity is defined as:
    $\small Sim_{p}(c_{p_{i}},c_{p_{j}})=\frac{|L_{so}|}{{d(c_{p_{i}})+d(c_{p_{j}})-|L_{so}|}},$ (6)
    where $|L_{so}|$ denotes the number of common subject-object associations and $d(c_{p_{i}})$ denotes the sample number with the predicate category $c_{p_{i}}$ in the dataset.
  • Context Similarity: This measure evaluates the extent to which two predicates appear in similar environmental contexts by comparing their co-occurring entities:
    $\small Sim_{c}(c_{p_{i}},c_{p_{j}})=\frac{|L_{e}|}{{d_{co}(c_{p_{i}})+d_{co}(c_{p_{j}})-|L_{e}|}},\$ (7)
    where $|L_{e}|$ denotes the number of entity categories that co-occur with both $c_{p_{i}}$ and $c_{p_{j}}$ in the same image. The term $d_{co}(c_{p_{i}})$ represents the number of images in which predicate category $c_{p_{i}}$ appears.
  • Semantic Similarity: To capture semantic alignment, we compute the Euclidean distance between two predicate categories in a pre-trained embedding space:
    $Sim_{s}(c_{p_{i}},c_{p_{j}})=||\bm{e}_{p_{i}}-\bm{e}_{p_{j}}||_{2}.$ (8)
The final similarity score for predicate clustering is a weighted sum of these three measures.
(2) Predicate Category Merging. Once predicate clusters are established, associated subject and object categories within each cluster are merged to ensure consistency. Overlapping subject categories among clustered predicates are unified, guaranteeing consistent treatment of interchangeable subjects across relational structures. Likewise, object categories governed by these clustered predicates are merged, preserving the relational and contextual properties of entity substitutions.
For instance, if the predicates standing on and sitting on belong to the same cluster, their subject categories (e.g., dog, cat, person) and object categories (e.g., chair, floor, bed) are merged accordingly.
(3) Entity Category Clustering. After merging, we further cluster entity categories within the same predicate clusters to identify interchangeable entities. The clustering process follows a methodology similar to predicate clustering and employs three similarity measures. Fig. 6 shows an example of pattern and context similarity between cat and dog:
  • Pattern Similarity: This metric evaluates the extent to which two entity categories participate in similar relational structures. It is computed as:
    $Sim_{p}(c_{e_{i}},c_{e_{j}})$ $=\frac{|L_{sp}|}{{d_{out}(c_{e_{i}})+d_{out}(c_{e_{j}})-|L_{sp}|}}$ (9)
    $+\frac{|L_{po}|}{{d_{in}(c_{e_{i}})+d_{in}(c_{e_{j}})-|L_{po}|}},$
    where $|L_{sp}|$ and $|L_{po}|$ represent the number of common subject-predicate and predicate-object associations. Terms $d_{in}(c_{e_{i}})$ and $d_{out}(c_{e_{i}})$ denote the number of incoming and outgoing edges of entity category $c_{e_{i}}$.
  • Context Similarity: Similarly, this similarity measures the context’s overlap of two entity categories in images:
    $\small Sim_{c}(c_{e_{i}},c_{e_{j}})=\frac{|L_{e}|}{{d_{co}(c_{e_{i}})+d_{co}(c_{e_{j}})-|L_{e}|}}.\$ (10)
  • Semantic Similarity: As with predicates, it’s the distance between two entity semantic embeddings:
    $Sim_{s}(c_{e_{i}},c_{e_{j}})=||\bm{e}_{e_{i}}-\bm{e}_{e_{j}}||_{2}.$ (11)
Entire clustering algorithm and results are in the appendix.
Figure 6: Illustration of the pattern and context similarity between entity categories cat and dog. a) White boxes are the behavior patterns common to two categories. b) Blue and red circles are entity categories that co-occur with cat and dog, respectively.
Entity Feature Selection and Replacement. This step selects an entity feature from the feature bank based on the selected entity category and replaces it with the query triplet. Specifically, we first update a new entity category of query triplet (e.g., cat-laying on-bed). Then, we select all triplet features with the same category as query triplet from feature bank as candidates (cf., Fig. 5(a)). And we utilize spatial restriction to filter unreasonable triplet features.
Spatial Restriction. In real augmentation, even if the two entity categories appear to be interchangeable, the positions of the subject and object after replacement may not make sense. For example, in VG dataset (cf.. appendix), though category hat and shoe are in the same cluster, directly replacing hat with shoe would place the shoe on the head and make it contrary to common sense.
To further ensure the rationality of the replacement of the two entity features, we use the cosine similarity of the relative spatial position of subject-object of the query triplet and the selected triplet from the feature bank as restriction:
[…]
where $\bm{p}_{i}$ is the spatial vector from the center of the subject bbox to the center of object bbox. When the ${Sim}_{d}$ between $\bm{p}_{i}$ of the query triplet and $\bm{p}_{j}$ of the selected triplet from the feature bank is larger than the threshold $\sigma$, the entity feature of the selected triplet is reasonable for replacement.
During replacement, we randomly select one of the reasonable entity features and replace it into query triplet and change the original entity target to the new entity category.
Highlights. In the original intrisic-CFA of CFA Li et al. (2023a), we clustered entities uniformly, ignoring predicate-driven semantic discrepancy (e.g., kite and umbrella are replaceable for holding, but an umbrella is ill-suited for flying in the sky where a kite fits). In this paper, we devise a relation-conditioned entity clustering, allowing more accurate, predicate-aware entity groupings. This approach, for instance, discerns the similarity of kite and umbrella for holding versus their distinct roles for flying in, yielding more precise entity selection.
3.2.2 Extrinsic-CFA
Extrinsic-CFA aims to enrich the feature of the query triplet through the extrinsic information (i.e., the context formed by all the entities of an image) of context triplets. Different from the intrinsic-CFA, we select context triplet during training for computation efficiency. Extrinsic-CFA includes three steps: context triplet selection, query triplet feature selection, and adaptive feature augmentation (cf., Fig. 5(b)). Then, we detail each step of extrinsic-CFA.
Context Triplet Selection. For each image, this step selects context triplets for extrinsic-CFA. Specifically, we randomly select context triplets with foreground or background predicate categories to provide more extrinsic information. For foreground context triplets, considering the large variation in the number of triplet samples for different predicate categories, we randomly sample the context triplets by using probability $p=(\eta-{\eta_{r}})/\eta\times\gamma$ during training, where $\gamma$ is hyperparameter. For background context triplets, we randomly select them with the same subject-object pair as the tail predicate triplet in the whole dataset.
Query Triplet Feature Selection. This step selects a query triplet with its features for the feature augmentation. Specifically, we select triplets with the same subject-object categories (e.g., pillow-bed) as the context triplet (e.g., pillow-on-bed) from feature bank as the candidate query triplets. Then, we implement the same spatial restriction as intrinsic-CFA to filter unreasonable triplets. Finally, we randomly select one of the reasonable triplets as the query triplet (e.g., pillow-laying on-bed), and its features are utilized for augmentation (cf.. Fig. 5(b)).
Adaptive Feature Augmentation. To enhance the extrinsic features of tail predicate triplets while concurrently mitigating the impact on other triplets of the original image, we devise a novel adaptive mixup operation. This operation blends a selected query triplet (e.g., pillow-laying on-bed) with a context triplet (e.g., pillow-on-bed), as depicted in Fig. 5(b). The adaptive mixup is formulated as follows:
[…]
$\tilde{\bm{v}}_{s}=\theta\bm{v}_{s}+(1-\theta){{\bm{v}}_{s}^{\prime}},$
$\tilde{\bm{v}}_{o}=\theta{\bm{v}}_{o}+(1-\theta){{\bm{v}}_{o}^{\prime}},$
$\tilde{\bm{u}}=\theta\bm{u}+(1-\theta){{\bm{u}}^{\prime}},$
where $\bm{e}_{p}$ and $\bm{e}_{p}^{\prime}$ denote the semantic embedding of relation category in context triplet (e.g., on) and selected query triplet (e.g., laying on), respectively. The parameters $\alpha$ is used to control the value range of $\theta$. $\bm{v}_{s}^{\prime}$ ($\bm{v}_{s}$) and $\bm{v}_{o}^{\prime}$ ($\bm{v}_{o}$) denote the visual features of the subject and object of the query triplet (context triplet), respectively. The $\bm{u}^{\prime}$ ($\bm{u}$) denotes the visual feature of the union box of the query triplet (context triplet). Similarly, the predicate target ground-truth $r$ of the context triplet is mixed:
$\tilde{r}=\theta{r}+(1-\theta){r^{\prime}},$ (14)
where $\theta$ is utilized to control the degree of mixup operation.
After the mixup operation, the query triplet feature can be enhanced in context modeling (cf.. Fig. 4) by leveraging the extrinsic features of the context triplet.
Highlights. In CFA Li et al. (2023a), we employed a fixed $\theta$ for feature augmentation. It fails to adapt to varying semantic similarities between query and context triplets, risking under-enhancement of tail predicate triplets or feature corruption via suboptimal contextual blending. In this paper, we propose an adaptive feature augmentation strategy, enabling the mixup coefficient $\theta$ to dynamically adjust based on predicate similarity between query and context triplets. This facilitates targeted enrichment of tail predicates using relevant contexts while preventing feature distortion from less aligned ones, ultimately boosting model robustness.
3.3 Image-level CFA
Although the above mentioned methods effectively augment the diversity of both extrinsic and intrinsic features of predicates, the constructed feature space is still limited in the dataset domain. In order to maximally infer the real feature space of predicates and learn the true data distribution, we also propose an image-level augmentation method that leverages large generative models Li et al. (2023b) to create new visual appearance of relation triplets. Image-level CFA includes three steps: context triplet and query triplet selection, augmented triplet image generation, and quality estimation (cf., Fig. 5(c)). Then, we detail each step of Image-level CFA.
Context Triplet and Query Triplet Selection. For each image, this step selects context triplets and query triplet to be qugmented for image-level augmentation. We select context triplets with foreground predicate categories to provide extrinsic information. We apply the same sampling strategy as the way of extrinsic-CFA in Sec. 3.2.2, i.e., we randomly sample the context triplets by using probability $p=(\eta-{\eta_{r}})/\eta\times\gamma$ during training, where $\gamma$ is hyperparameter. We also apply the same strategy as the way of extrinsic-CFA in Sec. 3.2.2 to select query triplet, but we only use its categories and grounding information of entities and predicate. Then we send them with context triplets together to generate new visual images.
Augmented Triplet Image Generation. This step is to generate new image with context and query triplets. The main challenge is to ensure the rationality of generated image with query triplet. Taking the power of recent generative models, e.g., GLIGEN Li et al. (2023b), we generate new samples conditioned on more input information apart from text to control the generation process, e.g., bounding boxes and reference images.
Specifically, for each image $I$, suppose that the subject and object of the selected query triplet is $i$ and $j$, and the corresponding bounding boxes are $\bm{b}_{i},\bm{b}_{j}\in\mathcal{B}$ respectively. Our aim is to generate images with new visual appearance of entity $i$ and $j$ conditioned on the context information. We send the instruction $\bm{g}$ to the grounded text-to-image model GLIGEN Li et al. (2023b), which is a composition of reference image and query triplet information:
$\bm{g}=(\bm{e}_{i},\bm{e}_{j},r_{ij},I),\text{with}\;\bm{e}_{i}=(\bm{b}_{j},o_{i}\in\mathcal{O})$ (15)
where $\bm{e}_{i}$ are the information of entity $i$ with its category and bounding box, and $r_{ij}\in\mathcal{R}$ is the triplet predicate category. After pretrained with multiple input modalities besides text input, GLIGEN Li et al. (2023b) is able to generate improved quality image with grounding conditions. Some generated samples are displayed in Fig. 8.
Quality Estimation. After generating the new visual appearance of entities, we have to check their quality and make sure their category are remained consistent. We send the cropped regions of generated entities and triplets into the pretrained CLIP model Radford et al. (2021) to predict the classification scores. Only the correctly classified samples will be selected as the augmented datas to participate in the training process of SGG models.
3.4 Training Objectives
Cross-Entropy Losses. The optimization objective of commonly used SGG models mainly includes the cross-entropy of entity and relation classification, the loss functions are:
${L_{obj}}=\sum\nolimits_{i}{\mathtt{XE}({\widehat{o}_{i}},{o_{i}})},\;{L_{rel}}=\sum\nolimits_{ij}{\mathtt{XE}({\widehat{r}_{ij}},{r_{ij}})},$ (16)
where ${\widehat{o}_{i}}$ is the predicted entity category and ${o}_{i}$ is the ground-truth entity category. ${\widehat{r}_{ij}}$ is the predicted predicate category and ${r}_{ij}$ is the ground-truth predicate category.
Contrastive Loss. Due to the pattern changes between the entity features after mixup and original entity features, using only XE losses may result in performance drops in entity classification. To maintain the discriminative entity features after mixup, we further apply a contrastive loss Chen et al. (2020):
$L_{cl}=-\log\frac{{\exp(Si{m_{e}}({{\bm{z}}_{i}},{\bm{z}}_{j})/\tau)}}{{\sum\nolimits_{k=1}^{2M}{{1_{[k\neq i]}}\exp(Si{m_{e}}({{\bm{z}}_{i}},{{\bm{z}}_{k}})/\tau)}}},$ (17)
[…]
Training. During training, the total loss includes the cross-entropy losses $L_{obj}$, $L_{rel}$ and the extra contrastive loss $L_{cl}$:
$L_{total}=L_{rel}+L_{obj}+\beta L_{cl},$ (18)
where $\beta$ is used to regulate the magnitude of loss.
SGG Models PredCls SGCls SGGen
mR@K R@K Mean mR@K R@K Mean mR@K R@K Mean
50 100 50 100 50 100 50 100 50 100 50 100
Motifs  Zellers et al. (2018)${}_{\textit{CVPR'18}}$ 16.5 17.8 65.5 67.2 41.8 8.7 9.3 39.0 39.7 24.2 5.5 6.8 32.1 36.9 20.3
VCTree Tang et al. (2019)${}_{\textit{CVPR'19}}$ 17.1 18.4 65.9 67.5 42.2 10.8 11.5 45.6 46.5 28.6 7.2 8.4 32.0 36.2 20.9
Transformer Vaswani et al. (2017)${}_{\textit{NIPS'17}}$ 17.9 19.6 63.6 65.7 41.7 9.9 10.5 38.1 39.2 24.4 7.4 8.8 30.0 34.3 20.1
BGNN Li et al. (2021b)${}_{\textit{CVPR'21}}$ 30.4 32.9 59.2 61.3 45.9 14.3 16.5 37.4 38.5 26.7 10.7 12.6 31.0 35.8 22.5
Motifs+PCPL Yan et al. (2020)${}_{\textit{ACMMM'20}}$ 24.3 26.1 54.7 56.5 40.4 12.0 12.7 35.3 36.1 24.0 10.7 12.6 27.8 31.7 20.7
Motifs+DLFE Chiou et al. (2021)${}_{\textit{ACMMM'21}}$ 26.9 28.8 52.5 54.2 40.6 15.2 15.9 32.3 33.1 24.1 11.7 13.8 25.4 29.4 20.1
Motifs+BPL-SA Guo et al. (2021)${}_{\textit{ICCV'21}}$ 29.7 31.7 50.7 52.5 41.2 16.5 17.5 30.1 31.0 23.8 13.5 15.6 23.0 26.9 19.8
Motifs+NICE Li et al. (2022a)${}_{\textit{CVPR'22}}$ 29.9 32.3 55.1 57.2 43.6 16.6 17.9 33.1 34.0 25.4 12.2 14.4 27.8 31.8 21.6
Motifs+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.9 33.6 54.7 56.7 44.0 16.8 17.9 32.5 33.4 25.2 12.4 14.9 26.4 30.6 21.1
Motifs+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 35.7 38.2 54.1 56.6 46.2 17.0 18.4 34.9 36.1 26.6 13.2 15.5 27.4 31.8 22.0
Motifs+DPL Jeon et al. (2024)${}_{\textit{ECCV'24}}$ 33.7 37.4 54.4 56.3 45.5 18.5 20.1 32.6 33.8 26.3 13.0 15.6 24.5 28.7 20.5
Motifs+MCFA (ours) 35.8 38.3 54.3 56.8 46.3 17.9 19.4 35.5 36.7 27.4 14.0 16.3 27.1 31.6 22.3
VCTree+PCPL Yan et al. (2020)${}_{\textit{ACMMM'20}}$ 22.8 24.5 56.9 58.7 40.7 15.2 16.1 40.6 41.7 28.4 10.8 12.6 26.6 30.3 20.1
VCTree+DLFE Chiou et al. (2021)${}_{\textit{ACMMM'21}}$ 25.3 27.1 51.8 53.5 39.4 18.9 20.0 33.5 34.6 26.8 11.8 13.8 22.7 26.3 18.7
VCTree+BPL-SA Guo et al. (2021)${}_{\textit{ICCV'21}}$ 30.6 32.6 50.0 51.8 41.3 20.1 21.2 34.0 35.0 27.6 13.5 15.7 21.7 25.5 19.1
VCTree+NICE Li et al. (2022a)${}_{\textit{CVPR'22}}$ 30.7 33.0 55.0 56.9 43.9 19.9 21.3 37.8 39.0 29.5 11.9 14.1 27.0 30.8 21.0
VCTree+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.3 33.9 53.0 55.0 43.1 16.5 18.1 32.9 33.8 25.3 11.5 14.0 25.4 29.3 20.1
VCTree+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 34.5 37.2 54.7 57.5 46.0 19.1 20.8 42.4 43.5 31.5 13.1 15.5 27.1 31.2 21.7
VCTree+DPL Jeon et al. (2024)${}_{\textit{ECCV'24}}$ 35.3 37.9 54.0 55.8 45.8 19.2 21.3 32.6 33.9 26.8 13.2 15.9 27.0 31.4 21.9
VCTree+MCFA (ours) 35.4 38.2 54.0 56.9 46.1 21.0 23.3 41.8 42.9 32.3 13.4 15.8 27.2 31.6 22.0
Transformer+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.8 34.5 51.8 53.8 42.7 17.4 19.1 32.6 33.5 25.7 12.5 15.0 25.5 29.6 20.7
Transformer+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 30.1 33.7 59.2 61.5 46.1 15.7 17.2 36.3 37.3 26.6 12.3 14.6 27.7 32.1 21.7
Transformer+MCFA (ours) 32.4 34.4 58.9 61.0 46.7 16.4 17.6 36.7 37.9 27.2 12.6 15.1 27.4 32.0 21.8
Table 1: Performance (%) of the SOTA trade-off SGG models on VG Krishna et al. (2017). “Mean” is the average of mR@50/100 and R@50/100.
SGG Models PredCls SGCls SGGen
mR@20 mR@50 mR@100 mR@20 mR@50 mR@100 mR@20 mR@50 mR@100
Motifs+TDE Tang et al. (2020)${}_{\textit{CVPR'20}}$ 18.5 25.5 29.1 9.8 13.1 14.9 5.8 8.2 9.8
Motifs+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 20.9 26.4 29.0 12.1 14.9 16.1 7.9 10.4 11.8
Motifs+RTPB Chen et al. (2022)${}_{\textit{AAAI'22}}$ 28.8 35.3 37.7 16.3 20.0 21.0 9.7 13.1 15.5
Motifs+PPDL Li et al. (2022c)${}_{\textit{CVPR'22}}$ 27.9 32.2 33.3 15.8 17.5 18.2 9.2 11.4 13.5
Motifs+GCL Dong et al. (2022)${}_{\textit{CVPR'22}}$ 30.5 36.1 38.2 18.0 20.8 21.8 12.9 16.8 19.3
Motif+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 30.1 36.3 38.7 17.1 20.8 22.1 10.8 14.6 17.3
Motifs+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.5 39.9 43.0 17.3 20.9 22.4 11.2 15.3 18.1
Motifs+MCFA (ours) 31.8 40.3 43.5 17.7 21.3 22.7 11.9 15.8 18.6
VCTree+TDE Tang et al. (2020)${}_{\textit{CVPR'20}}$ 18.4 25.4 28.7 8.9 12.2 14.0 6.9 9.3 11.1
VCTree+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 22.0 27.6 29.7 15.4 18.8 19.9 7.8 10.4 12.1
VCTree+RTPB Chen et al. (2022)${}_{\textit{AAAI'22}}$ 27.3 33.4 35.6 20.6 24.5 25.8 9.6 12.8 15.1
VCTree+PPDL Li et al. (2022c)${}_{\textit{CVPR'22}}$ 29.7 33.3 33.8 20.3 21.8 22.4 9.1 11.3 13.3
VCTree+GCL Dong et al. (2022)${}_{\textit{CVPR'22}}$ 31.4 37.1 39.1 19.5 22.5 23.5 11.9 15.2 17.5
VCTree+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 31.0 36.9 39.2 20.5 25.0 26.8 10.1 13.7 16.3
VCTree+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.6 39.2 42.5 21.5 26.3 28.3 10.8 15.1 17.9
VCTree+MCFA (ours) 32.3 40.2 43.2 21.7 27.2 29.2 11.0 15.3 18.0
Transformer+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 22.9 28.4 31.0 13.0 15.7 16.7 7.9 11.1 12.7
Transformer+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 27.4 33.3 35.9 15.7 19.1 20.4 11.4 15.0 17.7
Transformer+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.2 38.6 41.5 17.2 20.9 22.7 10.6 15.0 17.9
Transformer+MCFA (ours) 31.4 38.9 42.2 17.5 21.5 23.1 10.8 15.1 18.1
[…]
Models PredCls SGCls SGGen
mR@50/100 mR@50/100 mR@50/100
Motifs Zellers et al. (2018) 13.9 / 14.7 7.2 / 7.5 5.5 / 6.6
+CFA 31.7 / 33.8 14.2 / 15.2 11.6 / 13.2
+DPL Jeon et al. (2024) 31.6 / 33.9 13.3 / 14.4 11.1 / 13.1
+MCFA 32.5 / 33.9 14.7 / 15.8 12.0 / 13.8
VCTree Tang et al. (2019) 14.4 / 15.3 6.1 / 6.6 5.8 / 6.0
+CFA 33.4 / 35.1 14.1 / 15.0 10.8 / 12.6
+DPL Jeon et al. (2024) 31.4 / 34.3 13.3 / 14.9 11.9 / 13.7
+MCFA 33.8 / 35.5 14.3 / 15.4 11.1 / 13.1
Transformer Vaswani et al. (2017) 15.2 /16.1 7.5 / 7.9 6.9 / 7.8
+CFA 27.8 / 29.4 16.2 / 16.9 13.4 / 15.3
+MCFA 28.6 / 30.4 16.6 / 17.2 13.5 / 15.6
Table 3: Performance (%) of the SGG models on GQA Hudson and Manning (2019).
4 Experiments
4.1 Experimental Settings and Details
Tasks. We evaluated models in three tasks Xu et al. (2017): 1) Predicate Classification (PredCls): Predicting the predicate category given all ground-truth entity bboxes and categories. 2) Scene Graph Classification (SGCls): Predicting categories of the predicate and entity given all ground-truth entity bboxes. 3) Scene Graph Generation (SGGen): Detecting all entities and their pairwise predicates.
Metrics. We evaluated SGG models on three metrics: 1) Recall@K (R@K): It indicates the proportion of ground-truths that appear among the top-$K$ confident predicted relation triplets. 2) mean Recall@K (mR@K): It is the average of R@K scores which are calculated for each predicate category separately. 3) Mean: It is the average of all R@K and mR@K scores. Since R@K favors head predicates while mR@K favors tail predicates, the Mean can better reflect the overall performance of all predicates Li et al. (2022a).
Datasets. We conducted all experiments on two datasets: VG and GQA. 1) VG: We followed the widely-used strategy Xu et al. (2017) to split dataset (i.e., 75K/32K images for train/test, 150 object categories and 50 predicate categories). Besides, 5K images within the training set are sampled as val set following Tang et al. (2020). All predicate categories are divided into three groups {head, body, tail} based on the number of samples as Liu et al. (2019). 2) GQA: We followed the prior work Dong et al. (2022) to split dataset (i.e., 70%/30% of the images for train/test, 200 object categories and 100 predicate categories). Similarly, 5K images within the training set are sampled as val set. We utilized the same group category ratio as VG to divide all predicate categories into {head, body, tail} groups.
Implementation Details. Following the training protocol in prior SGG works Tang et al. (2020), we adopted the object detector Faster R-CNN with the ResNeXt-101-FPN Lin et al. (2017) backbone trained by Tang et al. (2020) to detect all the bounding boxes and extract their visual features. The parameters of the backbone were kept frozen during the training. This detector could achieve 28.14 mAP on the VG test set (i.e., using 0.5 IoU threshold for evaluation). To avoid messy hyperparamter tuning, most of the hyperparamters follow preivous works. More specifically, the hyperparameters $\lambda$ and $\gamma$ were set to 0.07 and 0.7 in MCFA (cf., Sec. 3.2). The weights of pattern, context and semantic similarity were 1.0, 1.0 and 0.01 in intrinsic-CFA (cf., Sec. 3.2.1). The hyperparameter $\alpha$ used to control the value range of adaptive mixup coefficient $\theta$ was set to 0.5 in extrinsic-CFA (cf., Sec. 3.2.2) The threshold $\sigma$ was set to 0.5 in spatial restriction (cf., Sec. 3.2.2). The $\beta$ was set to 0.1 to regulate the loss during training (cf., Sec. 3.4). In this paper, SGD optimizer was used to train the model. The batch size was set to 12 and the initial learning rate was set to 0.01. After the performance on the val set reached the plateau period, the learning rate would be decayed by 10 for two times. All experiments were carried out with PyTorch and NVIDIA 2080Ti GPU.
4.2 Comparison with State-of-the-Arts
Setting. Due to the model-agnostic nature, we equipped our MCFA with three strong two-stage SGG baselines: Motifs Zellers et al. (2018), VCTree Tang et al. (2019) and Transformer Vaswani et al. (2017), and they are denoted as Motifs+MCFA, VCTree+MCFA, and Transformer+MCFA, respectively. In addition, due to the limited diversity of tail predicate components Li et al. (2022d), it has a high correlation with the category of subject & object. Thus, we further equip our three models with the component prior knowledge collected from the dataset to further improve the performance of the tail predicates. And they are denoted as Motifs+MCFA, VCTree+MCFA, and Transformer+MCFA (More details about the priors are discussed in appendix).
Baselines. We compared our methods with the state-of-the-art (SOTA) models in the VG dataset (Table 1 and Table 2) and the GQA dataset (Table 3). Specifically, these models can be divided into three groups: 1) Model-specific designs: Motifs, VCTree, Transformer, and BGNN Li et al. (2021b). 2) Model-agnostic trade-off methods: consider the performance of all predicates comprehensively (i.e., higher Mean), e.g., PCPL Yan et al. (2020), DLFE Chiou et al. (2021), BP-LSA Guo et al. (2021), NICE Li et al. (2022a), IETrans Zhang et al. (2022), CFA Li et al. (2023a), and DPL Jeon et al. (2024). 3) Model-agnostic tail-focused methods: improve tail predicates performance at the expense of excessively sacrificing the head (i.e., higher mR@K and R@50 less than 50.0% on PredCls), e.g., TDE Tang et al. (2020), CogTree Yu et al. (2021), RTPB Chen et al. (2022), PPDL Li et al. (2022c), GCL Dong et al. (2022), and HML Deng et al. (2022). For a fair comparison, we compared with the methods in the last two groups.
Figure 7: The examples of reasonable entities for the query triplet in intrinsic-CFA. The lefts are original entities of the query triplets. The rights are some alternate entity categories (red) and their samples (gray boxes) for each query triplet to replace.
Figure 8: The examples of generated triplets for image-level augmentation.
Quantitative Results on VG. Analysis of the trade-off methods presented in Table 1 highlights two key observations: 1) Compared to three strong baselines (i.e., Motifs, VCTree, and Transformer), MCFA significantly enhances performance on the mR@K metric across all three experimental settings. 2) MCFA strikes the most effective trade-off between R@K and mR@K, evidenced by the highest Mean score. Notably, it surpasses the previous SOTA trade-off method, CFA Li et al. (2023a), which itself adeptly minimized performance decline on head predicates (R@K) while preserving tail performance (mR@K), in this crucial Mean metric under all settings. Turning to the tail-focused methods detailed in Table 2, MCFA, which incorporates further strategies dedicated to improving tail predicate performance, achieves the highest mR@K values. It outperforms prominent SOTA tail-focused approaches, i.e., HML Deng et al. (2022) and CFA Li et al. (2023a), on the mR@K metric across all evaluated settings. Additional experimental analyses are provided in the appendix.
Quantitative Results on GQA. As demonstrated in Table 3, MCFA also markedly enhances mR@K performance over three strong baselines (i.e., Motif, VCTree, and Transformer) on the large-scale GQA dataset. These results certify the generalizability and effectiveness of our method.
4.3 Ablation Studies
Component SGCls
IN EX-fg EX-bg mR@50 / 100 R@50 / 100 Mean
8.7 / 9.3 39.0 / 39.7 24.2
11.6 / 12.4 37.4 / 38.6 25.0
15.6 / 16.5 34.9 / 36.2 25.8
14.3 / 15.7 35.9 / 37.0 25.7
17.0 / 18.0 35.8 / 36.9 26.9
17.1 / 18.3 35.4 / 37.0 27.0
17.8 / 18.8 34.4 / 35.6 26.7
17.9 / 19.4 35.5 / 36.7 27.4
Table 4: Ablation study on each component in feature-level augmentation Krishna et al. (2017). Motifs Zellers et al. (2018) is used in all ablations on VG Krishna et al. (2017). IN: Replace the intrinsic features. Ex-fg: Mix up the extrinsic features of foreground triplets. Ex-bg: Mix up the extrinsic features of background triplets.
Similarity SGCls
Pattern Context Semantic mR@50 / 100 R@50 / 100 Mean
16.1 / 17.3 34.4 / 35.8 25.9
17.1 / 18.3 36.1 / 37.3 27.2
18.1 / 19.1 33.5 / 34.8 26.4
16.9 / 18.2 35.0 / 36.3 26.6
17.4 / 18.5 35.7 / 36.8 27.1
16.5 / 17.6 37.0 / 38.2 27.3
17.9 / 19.4 35.5 / 36.7 27.4
Table 5: Ablation study on each similarity in clustering of intrinsic-MCFA.
Feature-Level Augmentation. We assessed the impact of each component within our feature-level CFA using the Motifs Zellers et al. (2018) baseline under the SGCls setting. Feature-level CFA comprises three primary components: replacing intrinsic features (IN), mixing extrinsic features of foreground triplets (EX-fg), and mixing extrinsic features of background triplets (EX-bg). The results, reported in Table 8, lead to the following observations: 1) Employing only the IN component yields a marginal improvement in mR@K (e.g., 2.9%–3.1% gains) while slightly degrading R@K performance (a 1.1%–1.6% loss). This is likely attributable to the limited diversity among tail predicate triplets, meaning that feature variety remains constrained even after intrinsic feature replacement. 2) Both EX-fg and EX-bg components substantially enhance mR@K and also achieve R@K scores relative to the baseline. For instance, mR@100 increases to 18.0% from the baseline’s 9.3%, and R@100 maintains 18.8% with these components. 3) Combining all three components (IN, EX-fg, and EX-bg) achieves the optimal trade-off, i.e., the highest Mean score.
Similarity of Clustering in Intrinsic-CFA. We investigated the impact of three distinct similarity measures (pattern, context, and semantic) on clustering within intrinsic-CFA, utilizing the Motifs Zellers et al. (2018) model as a baseline under the SGCls setting. Results presented in Table 5 demonstrate that relying solely on pattern similarity yields the poorest SGG performance. This deficiency occurs since pattern-based similarity alone disregards crucial contextual and semantic information, which is essential for guiding meaningful cluster formation. For instance, while both a boat and a car may share a “ridable” pattern, substituting a car for a boat would be semantically invalid in a maritime context (e.g., on the sea), as a car cannot operate there. Conversely, when contextual and semantic similarities are integrated into the analysis, SGG performance becomes substantially more robust, culminating in the highest Mean.
$K_{p}$ SGCls
mR@50 / 100 R@50 / 100 Mean
10 17.5 / 18.9 35.3 / 37.1 27.2
20 17.9 / 19.4 35.5 / 36.7 27.4
50 17.3 / 18.4 35.5 / 37.3 27.1
Table 6: Ablation study on the number of predicate clusters $K_{p}$ of intrinsic-CFA.
Different Cluster Number. We performed $K_{p}$in {10, 20, 50} to evaluate the impact of the number of predicate clusters under SGCls setting with Motifs Zellers et al. (2018). As illustrated in Li et al. (2023a), since when entity cluster number $K_{e}$ set to 15, there is a larger selection range of replaceable entity categories and richer feature diversity of tail predicate triplets, we use same entity cluster number setting. All results are reported in Table 6. We find that $K_{p}$ = 20 works best in mR@K. The reason may be that when $K_{p}$ is smaller, such as 10, the predicate clusters might be too coarse, grouping semantically distinct predicates together. This over-generalization can lead to less precise entity selection by potentially merging entities that play truly different roles across these distinct predicates, thereby reducing the discriminativeness for specific relations. When $K_{p}$ is larger (e.g., 50), the predicate clusters may become overly granular, which could prevent the model from capturing broader semantic similarities between closely related predicates. Consequently, this limits the effectiveness of relation-conditioned entity merging as each predicate cluster becomes too specific and offers fewer possibilities for generalization.
$\theta$ SGCls
mR@50 / 100 R@50 / 100 Mean
0.0 16.9 / 17.9 35.2 / 36.5 26.6
0.5 17.1 / 18.6 36.0 / 37.2 27.2
1.0 11.6 / 12.4 37.4 / 38.6 25.0
Adap 17.9 / 19.4 35.5 / 36.7 27.4
Table 7: Ablation study on mixup parameter $\theta$ of extrinsic CFA. Motifs Zellers et al. (2018) is used in all ablation on VG Krishna et al. (2017).
Mixup Parameter $\theta$ in Extrinsic-CFA. As mentioned in Sec. 3.2.2, $\theta$ indicates the proportion of selected features in final augmented features. We investigated $\theta$in {0.0, 0.5, 1.0, adap} under the SGCls setting with Motifs Zellers et al. (2018) in Table 7. When $\theta$ is too small, the query triplet has too much impact on the other triplets in the image of the context triplet, and when $\theta$ is too large, feature augmentation of the query triplet is not strong enough. Our adaptive mixup operation alleviates this by dynamically calibrating $\theta$ based on the semantic similarity between the query and context predicates. Such calibration facilitates the targeted enrichment of tail predicates using relevant contexts, while concurrently preventing feature distortion from less pertinent ones, ultimately enhancing robustness.
Effectiveness of Each Enhanced Strategy in MCFA. We verified the effectiveness of each enhanced strategy on the original CFA, as presented in Table 8. Key observations are as follows: 1) Employing relation-conditioned entity clustering effectively identifies more suitable entities, leading to an improvement in Recall while maintaining a comparable mean Recall. 2) The adaptive mixup operation enables tail instances to better leverage contextual information from head instances, thereby significantly boosting mean Recall. 3) The adoption of image-level augmentation can simultaneously enhance recall and mean recall, which proves the rationality of generated triplets. 4) Combining all strategies allows MCFA to achieve the best Mean performance.
Component SGCls
En-IN En-EX IMG mR@50 / 100 R@50 / 100 Mean
17.0 / 18.4 34.9 / 36.1 26.6
16.4 / 17.6 37.1 / 38.2 27.3
18.3 / 20.1 34.1 / 35.4 27.0
18.1 / 18.9 35.0 / 36.2 27.1
17.9 / 19.4 35.5 / 36.7 27.4
Table 8: Ablation study on each enhanced strategy in MCFA on the VG dataset Krishna et al. (2017). En-IN: Utilize relation-conditioned entity clustering in intrinsic-CFA. En-EX: Employ the adaptive feature augmentation in extrinsic-CFA. IMG: Use the image-level augmentation.
5 Qualitative Analysis
Visualization of Reasonable Entities in Intrinsic-CFA. Fig. 7 shows some alternative entities for query triplets. Although the visual features of these candidate entities can differ significantly from the original, they remain reasonable substitutions within the query triplet. For example, while zebra samples may vary considerably in color and texture from a giraffe, both are suitable because they share the common behavior of “eating plants”. Replacing a giraffe with a zebra in such instances introduces new visual information, thereby enriching the feature diversity associated with the eating predicate.
Visualization of Generated Triplets for Image-level Augmentation. Fig. 8 displays some generated triplets (cf., Sec. 3.3), illustrating our method’s capability to generate diverse examples (e.g., handle-on back of-toilet) and to utilize the original query triplet’s bounding box for their supervision.
6 Conclusion
In this paper, we revealed the drawbacks of existing re-balancing methods and discovered that the key challenge for unbiased SGG is to learn proper decision boundaries under the severe long-tailed predicate distribution. To address this, we proposed a model-agnostic MCFA framework to enrich the feature space of tail categories. It augments both intrinsic and extrinsic features of relation triplets from the feature-level and image-level. At the feature-level, a relation-conditioned entity clustering strategy was designed for intrinsic augmentation and an adaptive mix-up strategy for extrinsic augmentation, dynamically adjusting based on predicate similarity. For image-level augmentation, we devised a generative method leveraging powerful models like GLIGEN to create new visual appearances for triplets, thereby overcoming in-domain data limitations. Comprehensive experiments on the challenging VG and GQA datasets consistently demonstrated that MCFA’s effectiveness for unbiased SGG.
Data Availability All experiments are conducted on publicly available datasets; see the references cited.
Appendix
This appendix is organized as follows:
  • The hierarchical clustering process and results mentioned in Sec. 3.2.1 are shown in Sec. A.
  • A more detailed discussion of prior knowledge mentioned in Sec. 4.2 is shown in Sec. B.
  • More experimental comparisons and analyses are shown in Sec. C
Input: Predicate category set $\mathcal{C}_{P}=\{c_{p_{i}}\mid i=1,2,\ldots,N_{p}\}$;
Entity category set $\mathcal{C}_{E}=\{c_{e_{i}}\mid i=1,2,\ldots,N_{e}\}$;
Predicate cluster similarity function $Sim_{P}$;
Entity cluster similarity function $Sim_{E}$;
Target number of predicate clusters $K_{p}$;
Target number of entity clusters $K_{e}$ within each predicate cluster;
Output: Set of all predicate clusters $\mathcal{S}^{p}=\{s^{p}_{i}|i=1,2,\cdots,K_{p}\}$
Set of all entity clusters $\mathcal{S}^{e}=\{s^{e}_{i}|i=1,2,\cdots,K_{e}\}$ within each predicate cluster $s^{p}_{i}$
/* --- Step 1: Predicate Clustering --- */
/* Initialize each predicate category as a cluster */
for $i=1,2,\cdots,N_{p}$ do
       $s_{i}^{p}=\{c_{p_{i}}\}$
/* Initialize the size of each cluster $l^{p}_{i}$ and the similarity matrix $A^{p}_{sim}$/
for $i=1,2,\cdots,N_{p}$ do
       $l^{p}_{i}=1$
      for $j=1,2,\cdots,N_{p}$ do
             $A_{sim}^{p}(i,j)=Sim_{P}(s^{p}_{i},s^{p}_{j})$
            $A_{sim}^{p}(j,i)=A^{p}_{sim}(i,j)$
/* Merge the two most similar predicate cluster until the number of clusters is smaller than $K_{p}$/
while $\text{LEN}(\mathcal{S}^{p})>K_{p}$ do
       $s^{p}_{i},s^{p}_{j}=\text{SELECT\_MAX}(A^{p}_{sim}(i,j)/(l^{p}_{i}+l^{p}_{j}))$
      $s^{p}_{i}=\text{MERGE}(s^{p}_{i},s^{p}_{j})$
      $\mathcal{S}^{p}=\mathcal{S}^{p}-{s^{p}_{j}}$
      $l^{p}_{i}=l^{p}_{i}+l^{p}_{j}+1$
      $\text{UPDATE}(A^{p}_{sim})$
/* --- Step 2: Predicate Merging --- */
for each predicate cluster $s^{p}_{i}\in\mathcal{S}^{p}$ do
       for $c_{e_{i}}\in C_{E}$ do
             if $c_{e_{i}}$ is in the same triplet category with $c_{p_{i}}\in s^{p}_{i}$ then
                   $C^{s^{p}_{i}}_{E}\leftarrow c_{e_{i}}$
/* --- Step 3: Entity Clustering within each Predicate Cluster --- */
for each predicate cluster $s^{p}_{i}\in\mathcal{S}^{p}$ do
       /* Initialize each entity category as a cluster */
      for $i=1,2,\cdots,|C^{s^{p}_{i}}_{E}|$ do
             $s_{i}^{e}=\{c_{e_{i}}\}$
      /* Initialize the size of each cluster $l^{e}_{i}$ and the similarity matrix $A^{e}_{sim}$/
      for $i=1,2,\cdots,N_{e}$ do
             $l^{e}_{i}=1$
            for $j=1,2,\cdots,N_{e}$ do
                   $A_{sim}^{e}(i,j)=Sim_{E}(s^{e}_{i},s^{e}_{j})$
                  $A_{sim}^{e}(j,i)=A^{e}_{sim}(i,j)$
      /* Merge the two most similar entity cluster until the number of clusters is smaller than $K_{e}$/
      while $\text{LEN}(\mathcal{S}^{e})>K_{e}$ do
             $s^{e}_{i},s^{e}_{j}=\text{SELECT\_MAX}(A^{e}_{sim}(i,j)/(l^{e}_{i}+l^{e}_{j}))$
            $s^{e}_{i}=\text{MERGE}(s^{e}_{i},s^{e}_{j})$
            $\mathcal{S}^{e}=\mathcal{S}^{e}-{s^{e}_{j}}$
            $l^{e}_{i}=l^{e}_{i}+l^{e}_{j}+1$
            $\text{UPDATE}(A^{e}_{sim})$
return $\mathcal{S}^{e}$ for each predicate cluster $s^{p}_{i}$;
Algorithm 1 Relation-Conditioned Entity Clustering
Figure 9: The results of relation clustering on VG Krishna et al. (2017).
Figure 10: The results of entity clustering on VG Krishna et al. (2017).
Appendix A Relation-Conditioned Entity Clustering
As mentioned in Sec. 3.2.1, we use relation-conditioned entity clustering to mine the replaceable entity categories. The specific implementation is detailed in Algorithm 1, which involves two main hierarchical clustering stages: first for predicates, and subsequently for entities, conditioned on these predicate clusters. Both stages utilize a similar iterative merging procedure.
The cluster similarity functions, $Sim_{P}$ for predicates and $Sim_{E}$ for entities, are the weighted sum of pattern, context, and semantic similarities. In practice, we treat clusters as a collection of multiple entity categories, and the pattern and context similarity can be transferred directly from class-level to cluster-level. For semantic similarity, we take the mean similarity for the categories between clusters.
After initializing the variables, the predicate and entity clustering stages share a similar logic in each iteration. First, we select the two most similar clusters (SELECT_MAX) based on similarity matrix ($A^{p}_{sim}$ for predicates or $A^{p}_{sim}$ for entities) and size of cluster ($l^{p}$ or $l^{e}$). Given that the distribution of different relation triplets in datasets is often highly imbalanced, we follow Wei et al. (2020) to employ the cluster size as a penalty term. This aims to prevent categories (whether predicate or entity) with a large proportion of samples from unduly dominating the clustering process. Subsequently, we merge the predicate or entity categories in cluster $s^{p}_{j}$ or $s^{e}_{j}$ into $s^{p}_{i}$ or $s^{e}_{i}$ (MERGE), remove cluster $s^{p}_{j}$ / $s^{e}_{j}$, and update the corresponding cluster size. Since the merged cluster has changed, we recalculate the similarity of other clusters to it and then update $A^{p}_{sim}$ / $A^{e}_{sim}$ (UPDATE). The above steps are repeated until the number of clusters is equal to $K^{p}$ for predicates or $K^{e}$ for entities.
The predicate clustering result and entity clustering results are shown in Fig. 9 and Fig. 10, respectively. Obviously, some predicate and entity categories that share common pattern are grouped into the same cluster, e.g., lying on and laying on (predicates), and dog and cat (entities, grouped within relevant predicate cluster contexts).
Figure 11: Performance(%) comparison between Motifs Zellers et al. (2018) and Motifs+CFA over all predicates on test set of VG Krishna et al. (2017). The orange area denotes the predicate distribution of training set.
Appendix B Component Prior Knowledge
As discussed in prior SGG works Li et al. (2022d), the statistic prior of the predicate distribution under a given condition can improve the mR@K performance of the unbiased SGG. In the inference phase, we calculate statistic component prior $b_{s,o,r}$, and add it to the predicted logits to predict predicate category:
${{{b}}_{s,o,r}}=-\log\frac{{{count_{s,o,r}}}}{{\sum\nolimits_{i=1}^{H}{{count_{s,o,i}}}}},$ (19)
where $H$ is the number of predicate categories. As for the prior “Triplet”, “Subject”, and “Object”, $count_{s,o,r}$ is the amount of the triplets whose predicate category is $r$ in the training set given subject-object pair, subject, and object respectively.
Appendix C Extra Comparisons and Analyses
Strategy PredCls
mR@50 / 100 R@50 / 100 Mean
Motifs Zellers et al. (2018) 16.5 / 17.8 65.6 / 67.2 41.8
+Reweight Tang et al. (2020) 30.8 ↑14.3 / 34.5 ↑16.7 36.1 / 40.4 35.5
+Resample Tang et al. (2020) 18.5 ↑2.0 / 20.0 ↑2.2 64.6 / 66.7 42.5
+MCFA 35.8 ${}_{{\color[rgb]{0.1725,0.7059,0.3921}\definecolor[named]{pgfstrokecolor}{rgb}{0.1725,0.7059,0.3921}\uparrow\textbf{18.7}}}$ / 38.3 ${}_{{\color[rgb]{0.1725,0.7059,0.3921}\definecolor[named]{pgfstrokecolor}{rgb}{0.1725,0.7059,0.3921}\uparrow\textbf{20.5}}}$ 54.3 / 56.8 46.3
Table 9: Performance (%) of re-balancing strategies and MCFA on VG Krishna et al. (2017).
Figure 12: The results of scene graphs generated by Motifs (blue) and Motifs-MCFA (green) on VG Krishna et al. (2017). Green predicates are correct (i.e., match GT), brown predicates are acceptable (i.e., does not match GT but still reasonable), and purple predicates are more informative (i.e., does not match GT and more reasonable).
Figure 13: The t-SNE visualization of randomly sampled instances of two predicates on the feature space before and after MCFA.
C.1 Comparison with Re-balancing Methods.
To demonstrate the superiority of MCFA compared with the prevalent re-balancing methods (i.e., reweight and resample) Tang et al. (2020), we conducted the three strategies on the baseline Motifs Zellers et al. (2018). The results under the PredCls setting are reported in Table 9. From the results, we can observe: 1) The Motifs baseline can achieve the best R@K. However, the high R@K is mainly due to the frequency bias of the dataset Tang et al. (2020), and they suffer severe drops in tail predicates. 2) The reweight method achieves better performance at mR@K, but it also sacrifices the performance of head predicates excessively, resulting in a low Mean. 3) The resample method keeps R@K high, but the improvement of mR@K is slight. The reason is that those decision boundaries may still be biased toward the head. 4) MCFA achieves the highest mR@K and maintains high R@K, i.e., it achieves the best trade-off over different predicate categories (highest performance on Mean).
C.2 Comparison over All Predicates
To further demonstrate the performance of each predicate, we displayed the predicate distribution of the training set in the VG dataset Krishna et al. (2017) and the performance of Motifs Zellers et al. (2018) and Motifs+MCFA on R@100 for each predicate category in Fig. 11. Obviously, our approach slightly compromises the performance of the head predicates (e.g., wearing), but greatly improves the tail predicates (e.g., flying in). This proves the superiority of our method in considering the performance of all predicates.
C.3 Qualitative Results.
Figure 12 shows some qualitative results generated by Motifs Zellers et al. (2018) and Motifs+MCFA under PredCls setting. We can observe that MCFA can not only predict more accurate predicates (e.g., on vsparked on and on vs.walking on), but also more fine-grained and informative predicates (e.g., holding vsusing).
C.4 Quality of Compositional Features.
We visualized the t-SNE distributions of features (from the previous layer of classification) in Fig. 13, demonstrating that our MCFA enhances the distinctiveness of tail predicates in feature space, i.e., it aids in finding suitable decision boundaries and does not harm to training, resulting in improved performance.

### v3

@@
-
+Compositional Feature Augmentation for Unbiased Scene Graph Generation
 Abstract
-Scene Graph Generation (SGG) aims to detect all the visual relation triplets $<$sub, pred, obj$>$ in a given image.
+Scene Graph Generation (SGG) aims to detect all the visual relation triplets <<sub, pred, obj>> in a given image.
 With the emergence of various advanced techniques for better utilizing both the intrinsic and extrinsic information in each relation triplet, SGG has achieved great progress over the recent years.
@@
 In this paper, we argue that all existing re-balancing strategies fail to increase the diversity of the relation triplet features of each predicate, which is critical for robust SGG.
-To this end, we propose a novel Multi-level Compositional Feature Augmentation (MCFA) strategy, which aims to mitigate the bias issue from the perspective of increasing the diversity of triplet features.
-Specifically, we enhance relationship diversity on not only feature-level, i.e., replacing the intrinsic or extrinsic visual features of triplets with other correlated samples to create novel feature compositions for tail predicates, but also image-level, i.e., manipulating the image to generate brand new visual appearance for triplets.
-Due to its model-agnostic nature, MCFA can be seamlessly incorporated into various SGG frameworks.
-Extensive ablations have shown that MCFA achieves a new state-of-the-art performance on the trade-off between different metrics.
-Keywords:
-Unbiased Scene Graph Generation Feature Augmentation Category ClusteringCompositional Learning
+To this end, we propose a novel Compositional Feature Augmentation (CFA) strategy, which is the first unbiased SGG work to mitigate the bias issue from the perspective of increasing the diversity of triplet features.
+Specifically, we first decompose each relation triplet feature into two components: intrinsic feature and extrinsic feature, which correspond to the intrinsic characteristics and extrinsic contexts of a relation triplet, respectively.
+Then, we design two different feature augmentation modules to enrich the feature diversity of original relation triplets by replacing or mixing up either their intrinsic or extrinsic features from other samples.
+Due to its model-agnostic nature, CFA can be seamlessly incorporated into various SGG frameworks.
+Extensive ablations have shown that CFA achieves a new state-of-the-art performance on the trade-off between different metrics.
+Corresponding author. Work was done when Lin Li visited HKUST.
 1 Introduction
-As one of the fundamental comprehensive visual scene understanding tasks, Scene Graph Generation (SGG) has attracted unprecedented interest from our community and has made great progress in recent years Lu et al.
-(2016); Zhang et al.
-(2017); Zellers et al.
-(2018); Chen et al.
-(2019a, b); Tang et al.
-(2020); Li et al.
-(2021b, 2022a).
+As one of the fundamental comprehensive visual scene understanding tasks, Scene Graph Generation (SGG) has attracted unprecedented interest from our community and has made great progress in recent years [29, 46, 44, 2, 3, 35, 24, 26, 21, 43, 23, 32].
 Specifically, SGG aims to transform an image into a visually-grounded graph representation (i.e., scene graph) where each node represents an object instance with a bounding box and each directed edge represents the corresponding predicate between the two objects.
-Thus, each scene graph can also be formulated as a set of visual relation triplets (i.e., $<$sub, pred, obj$>$).
+Thus, each scene graph can also be formulated as a set of visual relation triplets (i.e., <<sub, pred, obj>>).
 Since such structural representations can provide strong explainable potentials, SGG has been widely-used in various downstream tasks, such as visual question answering, image retrieval, and captioning.
@@
 The entity prediction is for the green box, and the predicate prediction is for the relation between the red and green boxes.
-(b) Illustration of the diversity of feature space and decision boundary between on and laying on before and after using re-balancing and MCFA.
+(b) Illustration of the diversity of feature space and decision boundary between on and laying on before and after using re-balancing and CFA.
 Each sample denotes the corresponding visual triplet features.
-In general, due to the extremely diverse visual appearance of different visual relation triplets, recent SGG methods all consider both intrinsic and extrinsic information for entity and predicate classification Zellers et al.
-(2018); Tang et al.
-(2019); Lu et al.
-(2021).
+In general, due to the extremely diverse visual appearance of different visual relation triplets, recent SGG methods all consider both intrinsic and extrinsic information for entity and predicate classification [44, 36, 30].
 By “intrinsic information”, we mean these intrinsic characteristics of the subjects and objects, such as their visual, semantic, and spatial features.
-For example in Fig.
-1(a), with the help of these intrinsic features, we can easily infer all possible entity categories (e.g., pole or surfboard) and predicate categories (e.g., on or holding) of each triplet.
+For example in Figure 1(a), with the help of these intrinsic features, we can easily infer all possible entity categories (e.g., pole or surfboard) and predicate categories (e.g., on or holding) of each triplet.
 However, sometimes it is still hard to confirm the exact correct predictions with only intrinsic information, especially for tiny objects.
 Thus, it is also essential to consider other “extrinsic information” in the same image, such as the context features from neighbor objects.
-As shown in Fig.
-1(a), after encoding the features of surrounding objects (e.g., wave and hand), we can easily infer that the categories of the entity and predicate should be surfboard and holding.
-Although numerous advanced techniques have been proposed to enhance SGG performance.
-Today’s SGG methods still fail to predict some informative predicates due to the ubiquitous long-tailed predicate distribution in prevalent SGG datasets Krishna et al.
-(2017).
-Such a distribution is characterized by a small amount categories with vast samples (head [^1] We directly use “tail”, “body”, and “head” categories to represent the predicate categories in the tail, body, and head parts of the number distributions of different predicates in SGG datasets, respectively.) and plenty of categories with rare samples (tail).
-As a result of the discrepancy of sample size among different categories, the learned decision boundary becomes improper (cf..
-Fig.
-1(b)), i.e., their predictions are biased towards the head predicates (e.g., on) and they are error-prone for the tail ones (e.g., laying on).
-To overcome the bias issue, the most prevalent unbiased SGG solutions are re-balancing strategies, e.g., sample re-sampling Li et al.
-(2021b) and loss re-weighting Yan et al.
-(2020); Chen et al.
-(2022); Li et al.
-(2022c); Lyu et al.
-(2022).
+As shown in Figure 1(a), after encoding the features of surrounding objects (e.g., wave and hand), we can easily infer that the categories of the entity and predicate should be surfboard and holding.
+Although numerous advanced techniques have been proposed to effectively leverage both intrinsic and extrinsic information, today’s SGG methods still fail to predict some informative predicates due to the ubiquitous long-tailed predicate distribution in prevalent SGG datasets [20].
+Such a distribution is characterized by few categories with vast samples (head [^1] We directly use “tail”, “body”, and “head” categories to represent the predicate categories in the tail, body, and head parts of the number distributions of different predicates in SGG datasets, respectively.) and many categories with rare samples (tail).
+Since the discrepancy of feature diversity and sample size among different categories, the learned decision boundary becomes improper (c.f.
+Figure 1(b)), i.e., their predictions are biased towards the head predicates (e.g., on) and they are error-prone for the tail ones (e.g., laying on).
+To overcome the bias issue, the most prevalent unbiased SGG solutions are re-balancing strategies, e.g., sample re-sampling [26, 49] and loss re-weighting [41, 1, 27, 31, 17].
 They alleviate the negative impact of long-tailed distribution by increasing samples or loss weights of tail classes.
 Then, the decision boundaries are adjusted to reduce the bias introduced by imbalanced distributions.
-However, we argue that all the existing re-balancing strategies only change the frequencies or contributions of existing relation triplets to improve the skewed decision boundaries, but fail to enhance the diversity of relation triplet features [^2] We use the “relation triplet feature” to represent the combination of both intrinsic and extrinsic features of each visual relation triplet.
-of each predicate, which is crucial to infer their complete feature space.
-This sparsity of feature diversity will lead to under-representation of tail categories, thereby making it hard to infer the complete data distribution, i.e., making it challenging to find the optimal direction to adjust the decision boundaries Chu et al.
-(2020); Vigneswaran et al.
-(2021).
-For example in Fig.
-1(b), the feature space of laying on is so sparse that the decision boundary can be adjusted within a large range.
+However, we argue that all the existing re-balancing strategies fail to increase the diversity of relation triplet features [^2] We use the “relation triplet feature” to represent the combination of both intrinsic and extrinsic features of each visual relation triplet.
+of each predicate, i.e., they only change the frequencies or contributions of existing relation triplet features (c.f.
+Figure 1(b)).
+Since these tail categories are under-represented, it is still hard to infer the complete data distribution, i.e., making it challenging to find the optimal direction to adjust the decision boundaries [6, 38].
+For example in Figure 1(b), the feature space of laying on is so sparse that the decision boundary can be adjusted within a large range.
 The performance of this naive adjustment without “complete” distribution is always sensitive to hyperparameters, i.e., excessively increasing the sample number or loss weight of tail predicates may cause some head predicate samples to be incorrectly predicted as tail classes (right solid line), and vice versa (left solid line).
 Figure 2: (a) Intrinsic-CFA: Replacing the entity feature of tail predicate triplet dog-laying on-bed from dog to cat to enhance the intrinsic feature.
-(b) Extrinsic-CFA: Mixing up the feature of tail predicate triplet pillow-laying on-bed into the context of pillow-on-bed to enhance the extrinsic feature.
-Therefore, in this paper, we focus on enhancing diversity of relation triplet features instead of re-balancing their frequencies to learn more robust feature space and decision boundaries for predicates.
-One intuitive idea is to borrow novel entity features or context features from other existing triplet samples to replace the original features, thereby constructing new triplet features for the target predicates.
-By deeply exploring the correlations of objects and predicates, we can reasonably select the features to replace, ensuring the reliability of the re-composed relation triplet features.
-However, this naive augmentation technique on feature-level manipulation has its inevitable limitations.
-On the one hand, directly replacing the original visual features may lead to unnatural transitions (e.g., fusing a pillow’s feature with another one that are inadvertently entangled with adjacent objects, such as bedside cupboard) as shown in Fig.
-3(b), further hurting the expressive ability of the model.
-On the other hand, the augmented samples are still limited to the in-domain distribution, and as a result, the trained models are unable to handle with out-of-domain predictions.
-Therefore, we also propose to manipulate relation triplet features on image-level with the help of powerful generative models as shown in Fig.
-3(c).
-Figure 3: (a) Context Image: The image which provides context information and is used to to augment with the triplet pillow-on-bed.
-(b) Feature-level Augmentation: Visualization of augmented results with the feature of tail predicate triplet pillow-laying on-bed selected from dataset.
-(c) Image-level Augmentation: Augmented new image with target triplet and context information by generative models.
-In this paper, we propose a novel Multi-level Compositional Feature Augmentation (MCFA) strategy for unbiased SGG, which tries to solve the bias issue by enhancing the diversity of relation triplet features on both feature-level and image-level, especially for tail predicates.
-As shown in Fig. 1(b), by enhancing the feature diversity of the tail predicates (e.g., laying on), SGG models can easily learn the proper decision boundaries (vs. re-balancing strategies).
-On feature-level, we apply different strategies for augmenting intrinsic and extrinsic feature diversity respectively.
-For intrinsic diversity, we replace the entity features (e.g., subject or object) of a tail predicate triplet with other “suitable” entity features.
-To determine the suitable entity categories for the augmentation, we propose a novel relation-conditioned entity clustering method to find the predicate-aware correlations between different entity categories, and then we regard the entity features from the same cluster are suitable.
+(b) Extrnisc-CFA: Mixing up the feature of tail predicate triplet pillow-laying on-bed into the context of pillow-on-bed to enhance the extrinsic feature.
+In this paper, we propose a novel Compositional Feature Augmentation (CFA) strategy for unbiased SGG, which tries to solve the bias issue by enhancing the diversity of relation triplet features, especially for tail predicates.
+Specifically, CFA consists of two components: an intrinsic feature augmentation (intrinsic-CFA) and an extrinsic feature augmentation (extrinsic-CFA), which enhance the intrinsic and extrinsic features2, respectively.
+As shown in Figure 1(b), by increasing the feature diversity of the tail predicates (e.g., laying on), SGG models can easily learn the proper decision boundaries (vs.
+re-balancing strategies).
+For intrinsic-CFA, we replace the entity features (e.g., subject or object) of a tail predicate triplet with other “suitable” entity features.
+To determine the suitable entity categories for the augmentation, we propose a new hierarchical clustering method to find the correlations between different entity categories, and then we regard the entity features from the same cluster are suitable.
 Specifically, we calculate the category correlation by pattern, context, and semantic similarities.
-For example in Fig.
-2 (a), the entity categories cat and dog are in the same cluster, i.e., we can augment the intrinsic feature by replacing the dog entity feature with a cat entity feature.
-For extrinsic diversity, we take advantage of the context or interactions of other triplets (i.e., context triplets) and enhance the features of tail predicate triplets by these context triplets.
+For example in Figure 2(a), the entity categories cat and dog are in the same cluster, i.e., we can augment the intrinsic feature by replacing the dog entity feature with a cat entity feature.
+For extrinsic-CFA, we take advantage of the context or interactions of other triplets (i.e., context triplets) and enhance the features of tail predicate triplets by these context triplets.
 Specifically, given a context triplet randomly selected from an image, we first select a reasonable tail predicate triplet as the target by limiting the categories and relative position of two objects.
-Then, to minimize the impact on the prediction of other triplets in the original image and make use of the extrinsic features of the context triplet, we use an adaptive mixup strategy to fuse the features of targeted tail predicate triplet into the context triplet and adaptively adjust mixup weights based on predicate semantic similarity.
-For example in Fig.
-2 (b), triplet pillow-laying on-bed is mixed up into the image of pillow-on-bed.
-To avoid introducing noise into augmented samples, we adaptively assign weights to the mixed triplets by considering the semantic similarity of relation triplets.
-On image-level, we aim to directly manipulate the images and generate brand-new visual appearances for target predicate triplets.
-A core challenge in this process is ensuring the semantic plausibility and spatial coherence of the generated images.
-To address this, we first employ a relative position constraint to filter only contextually appropriate images associated with target triplet categories for augmentation.
-For the synthesis process, we leverage the powerful grounded text-to-image generation capabilities of recent generative models (e.g., GLIGEN Li et al.
-(2023b)), as shown in Fig.
-3(c).
-Specifically, the model is conditioned on the original image, the target triplet’s categories, and corresponding entity grounding information to generate a new high-quality visual instance.
-Subsequently, we assess the quality of each generated image via a visual-language similarity score to ensure high fidelity and relevance.
-As the context triplets are selected from other predicate triplets and visual features of target triplets are novelly generated, both extrinsic and intrinsic diversity are enhanced under our image-level augmentation strategy.
-We evaluate MCFA on two most prevalent and challenging SGG datasets: Visual Genome (VGKrishna et al.
-(2017) and GQA Hudson and Manning (2019).
-Since MCFA is a model-agnostic debiasing strategy, it can be seamlessly incorporated into various SGG architectures and consistently improve their performance.
-Unsurprisingly, MCFA can achieve a new state-of-the-art performance on the trade-off between different metrics.
-Extensive ablations and results on multiple SGG tasks and backbones have shown the generalization ability and effectiveness of MCFA.
-This paper is a comprehensive and in-depth extension of our previous ICCV work Li et al. (2023a) with several significant improvements:
+Then, to minimize the impact on the prediction of other triplets in the original image and make use of the extrinsic features of the context triplet, we use a mixup operation to fuse the features of targeted tail predicate triplet into the context triplet.
+For example in Figure 2(b), triplet pillow-laying on-bed is mixed up into the image of pillow-on-bed.
+We evaluate CFA on two most prevalent and challenging SGG datasets: Visual Genome (VG[20] and GQA [16].
+Since CFA is a model-agnostic debiasing strategy, it can be seamlessly incorporated into various SGG architectures [^3] Following the mainstream and concurrent unbiased SGG works, we also only focus on two-stage frameworks.
+As for one-stage models, a similar idea can be applied at the image-level, and we leave it for future works.
+and consistently improve their performance.
+Unsurprisingly, CFA can achieve a new state-of-the-art performance on the trade-off between different metrics.
+Extensive ablations and results on multiple SGG tasks and backbones have shown the generalization ability and effectiveness of CFA.
+In summary, we make three contributions in this paper:
   1. 1.
-    We propose a novel relation-conditioned entity clustering strategy for feature-level intrinsic augmentation by leveraging the predicate-driven semantic knowledge, leading to more accurate groupings.
+    We reveal the issue of existing re-balancing methods, i.e., the lack of triplet feature diversity of tail categories.
+To this end, we are the first to tackle unbiased SGG from the perspective of increasing the diversity of triplet features.
   2. 2.
-    We propose an adaptive mix-up strategy for feature-level extrinsic augmentation, which dynamically adjust mix-up coefficient based on predicate similarity, effectively handling with the varying semantic context of triplets.
+    We propose the model-agnostic CFA for unbiased SGG, which is an efficient and novel compositional learning framework that spans the feature space of the tail categories by two independent plug-and-play modules.
   3. 3.
-    We devise a novel generative augmentation method to manipulate relation triplet features on the image-level and generate high-quality samples, thereby effectively boosting the performance of tail predicates.
-  4. 4.
-    Extensive qualitative and quantitative analyses demonstrate the effectiveness of our MCFA.
+    Extensive results show the effectiveness of CFA, i.e., it achieves a new SOTA performance on SGG benchmarks.
 2 Related Work
@@
 Biased predictions prevent further use of scene graphs in real-world applications.
-Recent unbiased SGG works can be divided into three main categories: 1) Re-balancing: It alleviates the negative impact of long-tailed predicate distribution by re-weighting or re-sampling Yan et al.
-(2020); Li et al.
-(2022c); Lyu et al.
-(2022); Li et al.
-(2021b); Desai et al.
-(2021).
-2) Unbiased Inference: It makes unbiased predictions based on biased models Tang et al.
-(2020); Yu et al.
-(2021).
-3) Noisy Label Learning: It reformulates SGG as a noisy label learning problem and corrects the noisy samples Li et al.
-(2022a, b).
-This work points out the drawbacks of existing re-balancing methods and studies unbiased SGG from the new perspective of multi-level feature augmentation.
+Recent unbiased SGG works can be roughly divided into three main categories: 1) Re-balancing: It alleviates the negative impact of long-tailed predicate distribution by re-weighting or re-sampling [41, 27, 31, 26, 8].
+2) Unbiased Inference: It makes unbiased predictions based on biased models [35, 42].
+3) Noisy Label Learning: It reformulates SGG as a noisy label learning problem and corrects these noisy samples [21, 22].
+In this work, we point out the drawbacks of existing re-balancing methods and study unbiased SGG from the new perspective of feature augmentation.
 Feature Augmentation.
 Data augmentation is a prevalent training trick to improve models’ performance.
-Conventional data augmentation methods DeVries and Taylor (2017); Zhong et al.
-(2020) usually synthesize new samples in a hand-crafted manner.
-Compared to these image-level data augmentation methods, feature augmentation is another efficient way to improve models’ generalizability by directly synthesizing samples in the feature space Chu et al.
-(2020); Li et al.
-(2021a).
+Conventional data augmentation methods [9, 48] usually synthesize new samples in a hand-crafted manner.
+Compared to these image-level data augmentation methods, feature augmentation is another efficient way to improve models’ generalizability by directly synthesizing samples in the feature space [6, 25].
 Compared to existing data augmentation applications, SGG is a sophisticated task that involves intrinsic and extrinsic features.
-In this work, we propose MCFA to enrich the diversity of relation triplet features for debiasing from both feature-level and image-level.
+In this work, we propose CFA to enrich the diversity of relation triplet features for debiasing.
 Compositional Learning (CL).
 CL has been successfully applied to various computer vision tasks.
-As for visual scene understanding, some Human Object Interaction (HOI) detection works Hou et al.
-(2020); Kato et al.
-(2018); Hou et al.
-(2021) compose new interaction samples that significantly benefit both low-shot and zero-shot settings.
-Inspired by these successes, we adopt this compositional approach for unbiased SGG.
-Our method enhances triplet feature diversity by first decomposing representations into their intrinsic and extrinsic constituents, and then augmenting these two feature streams.
-Figure 4: The illustration of unbiased SGG framework with MCFA on feature-level augmentation.
-Figure 5: The pipeline of feature-level and image-level augmentation strategy.
-Feature-level strategy consists of two approaches to generate new compositions of triplet features tailored for intrinsic and extrinsic diversity enhancement respectively, i.e., (a) Intrinsic-CFA and (b) Extrinsic-CFA.
-While image-level strategy generates new visual appearance of relation triplet by generative models to ensure the augmentation quality, i.e.(c) Image-CFA.
-3 Multi-level Compositional Feature Augmentation
+As for visual scene understanding, some Human Object Interaction detection works [14, 18, 15] compose new interaction samples that significantly benefit both low-shot and zero-shot settings.
+To the best of our knowledge, only two unpublished SGG work [13, 19] also uses CL.
+Compared to [13], we have several key differences: 1) They aim to generate new representations which are close to the original triplet.
+Instead, we try to increase the diversity of triplet features.
+2) They only change the entities and ignore the extrinsic features.
+3) Their augmentation strategies are mainly based on the spatial locations or IoU of the entities.
+The second work [19] primarily addresses classification tasks.
+Our work differs from it in two aspects: 1) CFA increases the diversity of tail predicate features by leveraging both intrinsic and extrinsic information.
+2) we aim to improve robustness and performance under long-tailed predicate distributions.
+Figure 3: The illustration of unbiased SGG framework with CFA.
+Figure 4: The pipeline of Intrinsic CFA (a) and Extrinsic CFA (b).
+The blue and green boxes represent operations on the query triplet features in Intrinsic-CFA and context triplet features in Extrinsic-CFA, respectively.
+3 Approach
 Given an image $I$, a scene graph is formally represented as $\mathcal{G}$ = $\mathcal{N}$, $\mathcal{E}$}, where $\mathcal{N}$ and $\mathcal{E}$ denote the set of all objects and their pairwise visual relations, respectively.
@@
 3.1.
-Then, we elaborate the pipeline of MCFA, including feature-level and image-level feature augmentation strategy (cf., Fig.
-4) in Sec.
-3.2 and Sec.
-3.3, respectively.
-Finally, we detail the training objectives in Sec.
-3.4.
+Then, we detailedly introduce our CFA in Sec.
+3.2, including intrinsic-/extrinsic-CFA (c.f.
+Figure 3).
+Finally, we introduce the training objectives in Sec.
+3.3.
 3.1 Revisiting the Two-Stage SGG Baselines
-Since the mainstream SGG methods are two-stage models, we review the two-stage SGG framework here Zellers et al.
-(2018); Tang et al.
-(2019).
+Since the mainstream SGG methods are two-stage models, we review the two-stage SGG framework3 here [44, 36].
 A typical two-stage SGG model involves three steps: proposal generation, entity classification, and predicate classification.
@@
 This step aims to generate all bbox proposals $\mathcal{B}$.
-Given an image $I$, they first utilize an off-the-shelf object detector (e.g., Faster R-CNN Ren et al.
-(2015)) to detect all the proposals $\mathcal{B}$ and their visual features $\{\bm{v}_{i}\}$.
+Given an image $I$, they first utilize an off-the-shelf object detector (e.g., Faster R-CNN [34]) to detect all the proposals $\mathcal{B}$ and their visual features $\{\bm{v}_{i}\}$.
 Entity Classification $P(\mathcal{O}|\mathcal{B},I)$.
-This step mainly predicts the entity category of each $\bm{b}_{i}$in $\mathcal{B}$.
+This step mainly predicts the entity category of each $\bm{b}_{i}$ $\in$ $\mathcal{B}$.
 Given a visual feature $\bm{v}_{i}$ and proposal $\bm{b}_{i}$, they use an object context encoder $\mathtt{Enc}_{obj}$ to extract the contextual entity representation $\bm{f}_{i}$:
@@
 $\bm{\tilde{f}}_{i}=\mathtt{Enc}_{rel}(\bm{v}_{i}\oplus\bm{f}_{i}\oplus\bm{w}_{i}),$ (4)
-where $\bm{w}_{i}$ is the GloVe embedding Pennington et al.
-(2014) of predicted $\widehat{o}_{i}$.
-It is worth noting that both $\mathtt{Enc}_{cls}$ and $\mathtt{Enc}_{rel}$ often adopt a sequence model (e.g., Bi-LSTM Zellers et al.
-(2018), Tree-LSTM Tang et al.
-(2019), or Transformer Yu et al.
-(2021)) to better capture context.
+where $\bm{w}_{i}$ is the GloVe embedding [33] of predicted $\widehat{o}_{i}$.
+It is worth noting that both $\mathtt{Enc}_{cls}$ and $\mathtt{Enc}_{rel}$ often adopt a sequence model (e.g., Bi-LSTM [44], Tree-LSTM [36], or Transformer [42]) to better capture context.
 After relation feature encoding, they use a relation classifier $\mathtt{Cls}_{rel}$ to predict the relation $\widehat{r}_{ij}$ between any two subject-object pairs:
@@
 where $\circ$ denotes element-wise product, and $\bm{u}_{ij}$ denotes the visual feature of the union box $\bm{b}_{ij}$.
-3.2 Feature-level CFA
+3.2 CFA: Compositional Feature Augmentation
 In this paper, we treat a relation triplet feature2 in SGG as a combination of two components: intrinsic feature and extrinsic feature, which refer to the intrinsic information (i.e., the subject and object itself) and extrinsic information (i.e., the contextual objects and stuff), respectively.
-Correspondingly, our feature-level CFA consists of intrinsic-CFA and extrinsic-CFA to augment these two types of features.
+Correspondingly, our CFA consists of intrinsic-CFA and extrinsic-CFA to augment these two types of features.
 For ease of presentation, we call the “targeted tail predicate triplet whose feature to be enhanced” as “query triplet”, and the “relation triplet whose image is used to provide context” as “context triplet”.
 Besides, to facilitate the feature augmentation, we store all the visual features of all tail predicate triplets (i.e., visual features $\bm{v}$ of two entities and their union feature $\bm{u}$) in a feature bank before training.
-Besides, we adopt the repeat factor $\eta$ = $\max(1,\eta_{r})$ Gupta et al.
-(2019); Li et al.
-(2021b) to sample images to provide enough context triplet and query triplet for augmentation, where $\eta_{r}$ = $\sqrt{\lambda/{f_{r}}}$, $f_{r}$ is the frequency of predicate category $r$ on the entire dataset, and $\lambda$ is hyperparameter.
-The unbiased SGG pipeline with feature-level CFA is shown in Fig.
-4.
+Besides, we adopt the repeat factor $\eta$ = $\max(1,\eta_{r})$ [12, 26] to sample images to provide enough context triplet and query triplet for augmentation, where $\eta_{r}$ = $\sqrt{\lambda/{f_{r}}}$, $f_{r}$ is the frequency of predicate category $r$ on the entire dataset, and $\lambda$ is hyperparameter.
+The unbiased SGG pipeline with CFA is shown in Figure 3.
 3.2.1 Intrinsic-CFA
 Intrinsic-CFA enhances the feature of query triplet by replacing its visual features of entities (i.e., subject and object).
-It consists of two steps: entity category selection and entity feature selection and replacement (cf..
-Fig.
-5(a)).
+It consists of two steps: entity category selection and entity feature selection and replacement (c.f.
+Figure 4(a)).
 During training, we randomly select a query triplet from a batch of images, and randomly select one of its entity features as input.
-Then we put it into the Intrinsic-CFA module (cf..
-Fig.
-4).
+Then we put it into the Intrinsic-CFA module (c.f.
+Figure 3).
 Next, we detailedly introduce each step.
 Entity Category Selection.
-This step is used for determining the category to which the entity feature of the query triplet is replaced.
-Firstly, we propose a novel relation-conditioned entity clustering that organizes entity categories based on their semantic role in a triplet.
-The framework consists of the following procedures.
-(1) Predicate Clustering. To maintain consistency in entity substitutions, we first group predicates into several clusters. The clustering process utilizes three similarity measures:
-  • Pattern Similarity: This metric quantifies the degree of functional overlap between predicates by analyzing the frequency with which they govern similar subject-object pairs.
-Formally, for two predicate categories $c_{p_{i}}$ and $c_{p_{j}}$, pattern similarity is defined as:
-    $\small Sim_{p}(c_{p_{i}},c_{p_{j}})=\frac{|L_{so}|}{{d(c_{p_{i}})+d(c_{p_{j}})-|L_{so}|}},$ (6)
-    where $|L_{so}|$ denotes the number of common subject-object associations and $d(c_{p_{i}})$ denotes the sample number with the predicate category $c_{p_{i}}$ in the dataset.
-  • Context Similarity: This measure evaluates the extent to which two predicates appear in similar environmental contexts by comparing their co-occurring entities:
-    $\small Sim_{c}(c_{p_{i}},c_{p_{j}})=\frac{|L_{e}|}{{d_{co}(c_{p_{i}})+d_{co}(c_{p_{j}})-|L_{e}|}},\$ (7)
-    where $|L_{e}|$ denotes the number of entity categories that co-occur with both $c_{p_{i}}$ and $c_{p_{j}}$ in the same image.
-The term $d_{co}(c_{p_{i}})$ represents the number of images in which predicate category $c_{p_{i}}$ appears.
-  • Semantic Similarity: To capture semantic alignment, we compute the Euclidean distance between two predicate categories in a pre-trained embedding space:
-    $Sim_{s}(c_{p_{i}},c_{p_{j}})=||\bm{e}_{p_{i}}-\bm{e}_{p_{j}}||_{2}.$ (8)
-The final similarity score for predicate clustering is a weighted sum of these three measures.
-(2) Predicate Category Merging.
-Once predicate clusters are established, associated subject and object categories within each cluster are merged to ensure consistency.
-Overlapping subject categories among clustered predicates are unified, guaranteeing consistent treatment of interchangeable subjects across relational structures.
-Likewise, object categories governed by these clustered predicates are merged, preserving the relational and contextual properties of entity substitutions.
-For instance, if the predicates standing on and sitting on belong to the same cluster, their subject categories (e.g., dog, cat, person) and object categories (e.g., chair, floor, bed) are merged accordingly.
-(3) Entity Category Clustering.
-After merging, we further cluster entity categories within the same predicate clusters to identify interchangeable entities.
-The clustering process follows a methodology similar to predicate clustering and employs three similarity measures.
-Fig.
-6 shows an example of pattern and context similarity between cat and dog:
-  • Pattern Similarity: This metric evaluates the extent to which two entity categories participate in similar relational structures. It is computed as:
-    $Sim_{p}(c_{e_{i}},c_{e_{j}})$ $=\frac{|L_{sp}|}{{d_{out}(c_{e_{i}})+d_{out}(c_{e_{j}})-|L_{sp}|}}$ (9)
-    $+\frac{|L_{po}|}{{d_{in}(c_{e_{i}})+d_{in}(c_{e_{j}})-|L_{po}|}},$
-    where $|L_{sp}|$ and $|L_{po}|$ represent the number of common subject-predicate and predicate-object associations.
-Terms $d_{in}(c_{e_{i}})$ and $d_{out}(c_{e_{i}})$ denote the number of incoming and outgoing edges of entity category $c_{e_{i}}$.
-  • Context Similarity: Similarly, this similarity measures the context’s overlap of two entity categories in images:
-    $\small Sim_{c}(c_{e_{i}},c_{e_{j}})=\frac{|L_{e}|}{{d_{co}(c_{e_{i}})+d_{co}(c_{e_{j}})-|L_{e}|}}.\$ (10)
-  • Semantic Similarity: As with predicates, it’s the distance between two entity semantic embeddings:
-    $Sim_{s}(c_{e_{i}},c_{e_{j}})=||\bm{e}_{e_{i}}-\bm{e}_{e_{j}}||_{2}.$ (11)
-Entire clustering algorithm and results are in the appendix.
+This step is used for determining the category to which the entity feature of query triplet is replaced.
+Firstly, we propose a novel hierarchical clustering strategy to mine potentially fungible entity categories.
+Then we randomly select an entity category from the same cluster for the next entity feature selection.
+For the query triplet dog-laying on-bed in Figure 4, the category cat is selected from the same cluster as dog.
+The categories in the same cluster are common in the behavior patterns (e.g., they can be ridden), contexts (e.g., they can appear in a scene with a street) as well as semantics.
+Accordingly, three kinds of similarity are used to measure the common characters of two entity categories: pattern similarity, context similarity, and semantic similarity.
 Figure 5: Illustration of the pattern and context similarity between entity categories cat and dog.
@@
 b) Blue and red circles are entity categories that co-occur with cat and dog, respectively.
+(i) Pattern Similarity.
+It measures the overlap of the behavior patterns of two entity categories [39].
+Figure 5(a) visualizes the common pattern between cat and dog (e.g., they all can be attached by the paw).
+Pattern similarity between two entity categories $c_{i}$ and $c_{j}$ is defined as:
+$Si{m_{p}}({c_{i}},{c_{j}})$ $=\frac{{|{L_{s}}|}}{{{d_{out}}({c_{i}})+{d_{out}}({c_{j}})-|{L_{s}}|}}$ (6)
+$+\frac{{|{L_{o}}|}}{{{d_{in}}({c_{i}})+{d_{in}}({c_{j}})-|{L_{o}}|}},$
+where $|L_{s}|$ ($|L_{o}|$) is the number of common pred-obj (sub-pred) classes of the triplet whose sub (obj) class is $c_{i}$ or $c_{j}$.
+$d_{in}(c_{i})$ ($d_{out}(c_{i})$) is the number of incoming (outgoing) edges of the entity with class $c_{i}$ in the whole dataset.
+(ii) Context Similarity.
+It measures the overlap of other entity categories in images of the two entity categories.
+Both cat and dog can appear in the same scene with entity categories, e.g., room in Figure 5(b).
+It is defined as:
+$\small Si{m_{c}}({c_{i}},{c_{j}})=\frac{{|{L_{e}}|}}{{{d_{co}}({c_{i}})+{d_{co}}({c_{j}})-|{L_{e}}|}},\\$ (7)
+where $|L_{e}|$ denotes the number of entity category intersections that appear in the same image with entity category $c_{i}$ or with $c_{j}$.
+$d_{co}(c_{i})$ is the number of entity instances in the whole dataset that appear in the same image as category $c_{i}$.
+(iii) Semantic Similarity. It is the cosine similarity between two entity categories on semantic embedding [33] space:
+[…]
+The final similarity function $Sim({c_{i}},{c_{j}})$ used in clustering is the weighted sum of above three similarities. Complete clustering algorithm and results are in the appendix.
 Entity Feature Selection and Replacement.
@@
 Specifically, we first update a new entity category of query triplet (e.g., cat-laying on-bed).
-Then, we select all triplet features with the same category as query triplet from feature bank as candidates (cf., Fig.
-5(a)).
+Then, we select all triplet features with the same category as query triplet from feature bank as candidates (c.f.
+Figure 4(a)).
 And we utilize spatial restriction to filter unreasonable triplet features.
@@
 In real augmentation, even if the two entity categories appear to be interchangeable, the positions of the subject and object after replacement may not make sense.
-For example, in VG dataset (cf..
+For example, in VG dataset (c.f.
 appendix), though category hat and shoe are in the same cluster, directly replacing hat with shoe would place the shoe on the head and make it contrary to common sense.
-To further ensure the rationality of the replacement of the two entity features, we use the cosine similarity of the relative spatial position of subject-object of the query triplet and the selected triplet from the feature bank as restriction:
+To further ensure the rationality of the replacement of the two entity features, we use the cosine similarity of the relative spatial position of subject-object of the query triplet and the selected triplet from the feature bank as the restriction:
 […]
@@
 During replacement, we randomly select one of the reasonable entity features and replace it into query triplet and change the original entity target to the new entity category.
-Highlights.
-In the original intrisic-CFA of CFA Li et al.
-(2023a), we clustered entities uniformly, ignoring predicate-driven semantic discrepancy (e.g., kite and umbrella are replaceable for holding, but an umbrella is ill-suited for flying in the sky where a kite fits).
-In this paper, we devise a relation-conditioned entity clustering, allowing more accurate, predicate-aware entity groupings.
-This approach, for instance, discerns the similarity of kite and umbrella for holding versus their distinct roles for flying in, yielding more precise entity selection.
 3.2.2 Extrinsic-CFA
 Extrinsic-CFA aims to enrich the feature of the query triplet through the extrinsic information (i.e., the context formed by all the entities of an image) of context triplets.
-Different from the intrinsic-CFA, we select context triplet during training for computation efficiency.
-Extrinsic-CFA includes three steps: context triplet selection, query triplet feature selection, and adaptive feature augmentation (cf., Fig.
-5(b)).
-Then, we detail each step of extrinsic-CFA.
+Different from the Intrinsic-CFA, we first select context triplet during training for computation efficiency.
+Extrinsic-CFA includes three steps: context triplet selection, query triplet feature selection, and feature augmentation (c.f.
+Figure 4(b)).
+Then, we will detailedly introduce each step of Extrinsic-CFA.
 Context Triplet Selection.
@@
 Then, we implement the same spatial restriction as intrinsic-CFA to filter unreasonable triplets.
-Finally, we randomly select one of the reasonable triplets as the query triplet (e.g., pillow-laying on-bed), and its features are utilized for augmentation (cf..
-Fig.
-5(b)).
-Adaptive Feature Augmentation.
-To enhance the extrinsic features of tail predicate triplets while concurrently mitigating the impact on other triplets of the original image, we devise a novel adaptive mixup operation.
-This operation blends a selected query triplet (e.g., pillow-laying on-bed) with a context triplet (e.g., pillow-on-bed), as depicted in Fig.
-5(b).
-The adaptive mixup is formulated as follows:
-[…]
-$\tilde{\bm{v}}_{s}=\theta\bm{v}_{s}+(1-\theta){{\bm{v}}_{s}^{\prime}},$
-$\tilde{\bm{v}}_{o}=\theta{\bm{v}}_{o}+(1-\theta){{\bm{v}}_{o}^{\prime}},$
-$\tilde{\bm{u}}=\theta\bm{u}+(1-\theta){{\bm{u}}^{\prime}},$
-where $\bm{e}_{p}$ and $\bm{e}_{p}^{\prime}$ denote the semantic embedding of relation category in context triplet (e.g., on) and selected query triplet (e.g., laying on), respectively.
-The parameters $\alpha$ is used to control the value range of $\theta$.
-$\bm{v}_{s}^{\prime}$ ($\bm{v}_{s}$) and $\bm{v}_{o}^{\prime}$ ($\bm{v}_{o}$) denote the visual features of the subject and object of the query triplet (context triplet), respectively.
-The $\bm{u}^{\prime}$ ($\bm{u}$) denotes the visual feature of the union box of the query triplet (context triplet).
-Similarly, the predicate target ground-truth $r$ of the context triplet is mixed:
+Finally, we randomly select one of the reasonable triplets as the query triplet (e.g., pillow-laying on-bed), and its features are utilized for augmentation (c.f.
+Figure 4(b)).
+Feature Augmentation.
+To reduce the impact on other triplets of the original image and enhance the extrinsic features of the tail predicate triplets, we perform mixup operation [47] between the selected query triplet (e.g., pillow-laying on-bed) and context triplet (e.g., pillow-on-bed), c.f.
+Figure 4(b).
+The mixup operation is written as:
+$\tilde{\bm{v}}_{s}=\theta\bm{v}_{s}+(1-\theta){{\bm{v}}_{s}^{\prime}},$ (10)
+$\tilde{\bm{v}}_{o}=\theta{\bm{v}}_{o}+(1-\theta){{\bm{v}}_{o}^{\prime}},$ (11)
+$\tilde{\bm{u}}=\theta\bm{u}+(1-\theta){{\bm{u}}^{\prime}},$ (12)
+where $\bm{v}_{s}^{\prime}$ ($\bm{v}_{s}$) and $\bm{v}_{o}^{\prime}$ ($\bm{v}_{o}$) denote the visual features of the subject and object of the query triplet (context triplet), respectively.
+The $\bm{u}^{\prime}$ ($\bm{u}$) denotes the visual feature of the uion box of the query triplet (context triplet).
+Similarly, predicate target ground-truth $r$ of the context triplet is mixed:
 $\tilde{r}=\theta{r}+(1-\theta){r^{\prime}},$ (13)
 where $\theta$ is utilized to control the degree of mixup operation.
-After the mixup operation, the query triplet feature can be enhanced in context modeling (cf.. Fig. 4) by leveraging the extrinsic features of the context triplet.
-Highlights.
-In CFA Li et al.
-(2023a), we employed a fixed $\theta$ for feature augmentation.
-It fails to adapt to varying semantic similarities between query and context triplets, risking under-enhancement of tail predicate triplets or feature corruption via suboptimal contextual blending.
-In this paper, we propose an adaptive feature augmentation strategy, enabling the mixup coefficient $\theta$ to dynamically adjust based on predicate similarity between query and context triplets.
-This facilitates targeted enrichment of tail predicates using relevant contexts while preventing feature distortion from less aligned ones, ultimately boosting model robustness.
-3.3 Image-level CFA
-Although the above mentioned methods effectively augment the diversity of both extrinsic and intrinsic features of predicates, the constructed feature space is still limited in the dataset domain.
-In order to maximally infer the real feature space of predicates and learn the true data distribution, we also propose an image-level augmentation method that leverages large generative models Li et al.
-(2023b) to create new visual appearance of relation triplets.
-Image-level CFA includes three steps: context triplet and query triplet selection, augmented triplet image generation, and quality estimation (cf., Fig.
-5(c)).
-Then, we detail each step of Image-level CFA.
-Context Triplet and Query Triplet Selection.
-For each image, this step selects context triplets and query triplet to be qugmented for image-level augmentation.
-We select context triplets with foreground predicate categories to provide extrinsic information.
-We apply the same sampling strategy as the way of extrinsic-CFA in Sec.
-3.2.2, i.e., we randomly sample the context triplets by using probability $p=(\eta-{\eta_{r}})/\eta\times\gamma$ during training, where $\gamma$ is hyperparameter.
-We also apply the same strategy as the way of extrinsic-CFA in Sec.
-3.2.2 to select query triplet, but we only use its categories and grounding information of entities and predicate.
-Then we send them with context triplets together to generate new visual images.
-Augmented Triplet Image Generation.
-This step is to generate new image with context and query triplets.
-The main challenge is to ensure the rationality of generated image with query triplet.
-Taking the power of recent generative models, e.g., GLIGEN Li et al.
-(2023b), we generate new samples conditioned on more input information apart from text to control the generation process, e.g., bounding boxes and reference images.
-Specifically, for each image $I$, suppose that the subject and object of the selected query triplet is $i$ and $j$, and the corresponding bounding boxes are $\bm{b}_{i},\bm{b}_{j}\in\mathcal{B}$ respectively.
-Our aim is to generate images with new visual appearance of entity $i$ and $j$ conditioned on the context information.
-We send the instruction $\bm{g}$ to the grounded text-to-image model GLIGEN Li et al.
-(2023b), which is a composition of reference image and query triplet information:
-$\bm{g}=(\bm{e}_{i},\bm{e}_{j},r_{ij},I),\text{with}\;\bm{e}_{i}=(\bm{b}_{j},o_{i}\in\mathcal{O})$ (15)
-where $\bm{e}_{i}$ are the information of entity $i$ with its category and bounding box, and $r_{ij}\in\mathcal{R}$ is the triplet predicate category.
-After pretrained with multiple input modalities besides text input, GLIGEN Li et al.
-(2023b) is able to generate improved quality image with grounding conditions.
-Some generated samples are displayed in Fig.
-8.
-Quality Estimation.
-After generating the new visual appearance of entities, we have to check their quality and make sure their category are remained consistent.
-We send the cropped regions of generated entities and triplets into the pretrained CLIP model Radford et al.
-(2021) to predict the classification scores.
-Only the correctly classified samples will be selected as the augmented datas to participate in the training process of SGG models.
+After the mixup operation, the query triplet feature can be enhanced in context modeling (c.f. Figure 3) by leveraging the extrinsic features of the context triplet.
 3.3 Training Objectives
@@
 Due to the pattern changes between the entity features after mixup and original entity features, using only XE losses may result in performance drops in entity classification.
-To maintain the discriminative entity features after mixup, we further apply a contrastive loss Chen et al.
-(2020):
+To maintain the discriminative entity features after mixup, we further apply a contrastive loss [4]:
 $L_{cl}=-\log\frac{{\exp(Si{m_{e}}({{\bm{z}}_{i}},{\bm{z}}_{j})/\tau)}}{{\sum\nolimits_{k=1}^{2M}{{1_{[k\neq i]}}\exp(Si{m_{e}}({{\bm{z}}_{i}},{{\bm{z}}_{k}})/\tau)}}},$ (15)
@@
 50 100 50 100 50 100 50 100 50 100 50 100
-Motifs  Zellers et al. (2018)${}_{\textit{CVPR'18}}$ 16.5 17.8 65.5 67.2 41.8 8.7 9.3 39.0 39.7 24.2 5.5 6.8 32.1 36.9 20.3
-VCTree Tang et al. (2019)${}_{\textit{CVPR'19}}$ 17.1 18.4 65.9 67.5 42.2 10.8 11.5 45.6 46.5 28.6 7.2 8.4 32.0 36.2 20.9
-Transformer Vaswani et al. (2017)${}_{\textit{NIPS'17}}$ 17.9 19.6 63.6 65.7 41.7 9.9 10.5 38.1 39.2 24.4 7.4 8.8 30.0 34.3 20.1
-BGNN Li et al. (2021b)${}_{\textit{CVPR'21}}$ 30.4 32.9 59.2 61.3 45.9 14.3 16.5 37.4 38.5 26.7 10.7 12.6 31.0 35.8 22.5
-Motifs+PCPL Yan et al. (2020)${}_{\textit{ACMMM'20}}$ 24.3 26.1 54.7 56.5 40.4 12.0 12.7 35.3 36.1 24.0 10.7 12.6 27.8 31.7 20.7
-Motifs+DLFE Chiou et al. (2021)${}_{\textit{ACMMM'21}}$ 26.9 28.8 52.5 54.2 40.6 15.2 15.9 32.3 33.1 24.1 11.7 13.8 25.4 29.4 20.1
-Motifs+BPL-SA Guo et al. (2021)${}_{\textit{ICCV'21}}$ 29.7 31.7 50.7 52.5 41.2 16.5 17.5 30.1 31.0 23.8 13.5 15.6 23.0 26.9 19.8
-Motifs+NICE Li et al. (2022a)${}_{\textit{CVPR'22}}$ 29.9 32.3 55.1 57.2 43.6 16.6 17.9 33.1 34.0 25.4 12.2 14.4 27.8 31.8 21.6
-Motifs+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.9 33.6 54.7 56.7 44.0 16.8 17.9 32.5 33.4 25.2 12.4 14.9 26.4 30.6 21.1
-Motifs+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 35.7 38.2 54.1 56.6 46.2 17.0 18.4 34.9 36.1 26.6 13.2 15.5 27.4 31.8 22.0
-Motifs+DPL Jeon et al. (2024)${}_{\textit{ECCV'24}}$ 33.7 37.4 54.4 56.3 45.5 18.5 20.1 32.6 33.8 26.3 13.0 15.6 24.5 28.7 20.5
-Motifs+MCFA (ours) 35.8 38.3 54.3 56.8 46.3 17.9 19.4 35.5 36.7 27.4 14.0 16.3 27.1 31.6 22.3
-VCTree+PCPL Yan et al. (2020)${}_{\textit{ACMMM'20}}$ 22.8 24.5 56.9 58.7 40.7 15.2 16.1 40.6 41.7 28.4 10.8 12.6 26.6 30.3 20.1
-VCTree+DLFE Chiou et al. (2021)${}_{\textit{ACMMM'21}}$ 25.3 27.1 51.8 53.5 39.4 18.9 20.0 33.5 34.6 26.8 11.8 13.8 22.7 26.3 18.7
-VCTree+BPL-SA Guo et al. (2021)${}_{\textit{ICCV'21}}$ 30.6 32.6 50.0 51.8 41.3 20.1 21.2 34.0 35.0 27.6 13.5 15.7 21.7 25.5 19.1
-VCTree+NICE Li et al. (2022a)${}_{\textit{CVPR'22}}$ 30.7 33.0 55.0 56.9 43.9 19.9 21.3 37.8 39.0 29.5 11.9 14.1 27.0 30.8 21.0
-VCTree+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.3 33.9 53.0 55.0 43.1 16.5 18.1 32.9 33.8 25.3 11.5 14.0 25.4 29.3 20.1
-VCTree+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 34.5 37.2 54.7 57.5 46.0 19.1 20.8 42.4 43.5 31.5 13.1 15.5 27.1 31.2 21.7
-VCTree+DPL Jeon et al. (2024)${}_{\textit{ECCV'24}}$ 35.3 37.9 54.0 55.8 45.8 19.2 21.3 32.6 33.9 26.8 13.2 15.9 27.0 31.4 21.9
-VCTree+MCFA (ours) 35.4 38.2 54.0 56.9 46.1 21.0 23.3 41.8 42.9 32.3 13.4 15.8 27.2 31.6 22.0
-Transformer+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.8 34.5 51.8 53.8 42.7 17.4 19.1 32.6 33.5 25.7 12.5 15.0 25.5 29.6 20.7
-Transformer+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 30.1 33.7 59.2 61.5 46.1 15.7 17.2 36.3 37.3 26.6 12.3 14.6 27.7 32.1 21.7
-Transformer+MCFA (ours) 32.4 34.4 58.9 61.0 46.7 16.4 17.6 36.7 37.9 27.2 12.6 15.1 27.4 32.0 21.8
-Table 1: Performance (%) of the SOTA trade-off SGG models on VG Krishna et al. (2017). “Mean” is the average of mR@50/100 and R@50/100.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Motifs+BPL-SA $1{}_{\textit{ICCV'21}}$ 29.7 31.7 50.7 52.5 41.2 16.5 17.5 30.1 31.0 23.8 13.5 15.6 23.0 26.9 19.8
+[…]
+[…]
+Motifs+CFA (ours) 35.7 38.2 54.1 56.6 46.2 17.0 18.4 34.9 36.1 26.6 13.2 15.5 27.4 31.8 22.0
+[…]
+[…]
+VCTree+BPL-SA $1{}_{\textit{ICCV'21}}$ 30.6 32.6 50.0 51.8 41.3 20.1 21.2 34.0 35.0 27.6 13.5 15.7 21.7 25.5 19.1
+[…]
+[…]
+VCTree+CFA (ours) 34.5 37.2 54.7 57.5 46.0 19.1 20.8 42.4 43.5 31.5 13.1 15.5 27.1 31.2 21.7
+[…]
+Transformer+CFA (ours) 30.1 33.7 59.2 61.5 46.1 15.7 17.2 36.3 37.3 26.6 12.3 14.6 27.7 32.1 21.7
+Table 1: Performance (%) of the SOTA trade-off SGG models on VG [20]. “Mean” is the average of mR@50/100 and R@50/100.
 SGG Models PredCls SGCls SGGen
 mR@20 mR@50 mR@100 mR@20 mR@50 mR@100 mR@20 mR@50 mR@100
-Motifs+TDE Tang et al. (2020)${}_{\textit{CVPR'20}}$ 18.5 25.5 29.1 9.8 13.1 14.9 5.8 8.2 9.8
-Motifs+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 20.9 26.4 29.0 12.1 14.9 16.1 7.9 10.4 11.8
-Motifs+RTPB Chen et al. (2022)${}_{\textit{AAAI'22}}$ 28.8 35.3 37.7 16.3 20.0 21.0 9.7 13.1 15.5
-Motifs+PPDL Li et al. (2022c)${}_{\textit{CVPR'22}}$ 27.9 32.2 33.3 15.8 17.5 18.2 9.2 11.4 13.5
-Motifs+GCL Dong et al. (2022)${}_{\textit{CVPR'22}}$ 30.5 36.1 38.2 18.0 20.8 21.8 12.9 16.8 19.3
-Motif+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 30.1 36.3 38.7 17.1 20.8 22.1 10.8 14.6 17.3
-Motifs+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.5 39.9 43.0 17.3 20.9 22.4 11.2 15.3 18.1
-Motifs+MCFA (ours) 31.8 40.3 43.5 17.7 21.3 22.7 11.9 15.8 18.6
-VCTree+TDE Tang et al. (2020)${}_{\textit{CVPR'20}}$ 18.4 25.4 28.7 8.9 12.2 14.0 6.9 9.3 11.1
-VCTree+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 22.0 27.6 29.7 15.4 18.8 19.9 7.8 10.4 12.1
-VCTree+RTPB Chen et al. (2022)${}_{\textit{AAAI'22}}$ 27.3 33.4 35.6 20.6 24.5 25.8 9.6 12.8 15.1
-VCTree+PPDL Li et al. (2022c)${}_{\textit{CVPR'22}}$ 29.7 33.3 33.8 20.3 21.8 22.4 9.1 11.3 13.3
-VCTree+GCL Dong et al. (2022)${}_{\textit{CVPR'22}}$ 31.4 37.1 39.1 19.5 22.5 23.5 11.9 15.2 17.5
-VCTree+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 31.0 36.9 39.2 20.5 25.0 26.8 10.1 13.7 16.3
-VCTree+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.6 39.2 42.5 21.5 26.3 28.3 10.8 15.1 17.9
-VCTree+MCFA (ours) 32.3 40.2 43.2 21.7 27.2 29.2 11.0 15.3 18.0
-Transformer+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 22.9 28.4 31.0 13.0 15.7 16.7 7.9 11.1 12.7
-Transformer+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 27.4 33.3 35.9 15.7 19.1 20.4 11.4 15.0 17.7
-Transformer+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.2 38.6 41.5 17.2 20.9 22.7 10.6 15.0 17.9
-Transformer+MCFA (ours) 31.4 38.9 42.2 17.5 21.5 23.1 10.8 15.1 18.1
 […]
+[…]
+[…]
+Motifs+PPDL $_{\textit{CVPR'22}} 27$.9 32.2 33.3 15.8 17.5 18.2 9.2 11.4 13.5
+[…]
+[…]
+Motifs+CFA (ours) 31.5 39.9 43.0 17.3 20.9 22.4 11.2 15.3 18.1
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+VCTree+CFA (ours) 31.6 39.2 42.5 21.5 26.3 28.3 10.8 15.1 17.9
+[…]
+[…]
+Transformer+CFA (ours) 31.2 38.6 41.5 17.2 20.9 22.7 10.6 15.0 17.9
+Table 2: Performance (%) of the SOTA tail-focused SGG models on VG [20]. $\ddagger$} means using the component prior knowledge.
 Models PredCls SGCls SGGen
 mR@50/100 mR@50/100 mR@50/100
-Motifs Zellers et al. (2018) 13.9 / 14.7 7.2 / 7.5 5.5 / 6.6
+Motifs [44] 13.9 / 14.7 7.2 / 7.5 5.5 / 6.6
 +CFA 31.7 / 33.8 14.2 / 15.2 11.6 / 13.2
-+DPL Jeon et al. (2024) 31.6 / 33.9 13.3 / 14.4 11.1 / 13.1
-+MCFA 32.5 / 33.9 14.7 / 15.8 12.0 / 13.8
-VCTree Tang et al. (2019) 14.4 / 15.3 6.1 / 6.6 5.8 / 6.0
+VCTree [36] 14.4 / 15.3 6.1 / 6.6 5.8 / 6.0
 +CFA 33.4 / 35.1 14.1 / 15.0 10.8 / 12.6
-+DPL Jeon et al. (2024) 31.4 / 34.3 13.3 / 14.9 11.9 / 13.7
-+MCFA 33.8 / 35.5 14.3 / 15.4 11.1 / 13.1
-Transformer Vaswani et al. (2017) 15.2 /16.1 7.5 / 7.9 6.9 / 7.8
+Transformer [37] 15.2 /16.1 7.5 / 7.9 6.9 / 7.8
 +CFA 27.8 / 29.4 16.2 / 16.9 13.4 / 15.3
-+MCFA 28.6 / 30.4 16.6 / 17.2 13.5 / 15.6
-Table 3: Performance (%) of the SGG models on GQA Hudson and Manning (2019).
+Table 3: Performance (%) of the SGG models on GQA [16].
 4 Experiments
 4.1 Experimental Settings and Details
+Figure 6: The examples of reasonable entities for the query triplet in intrinsic CFA.
+The lefts are original entities of the query triplets.
+The rights are some alternate entity categories (red) and their samples (gray boxes) for each query triplet to replace.
+Component PredCls
+IN EX-fg EX-bg mR@50 / 100 R@50 / 100 Mean
+16.5 / 17.8 65.6 / 67.2 41.8
+19.3 / 21.2 64.7 / 66.6 43.0
+25.6 / 27.8 63.0 / 64.8 45.3
+23.9 / 26.3 63.3 / 65.7 44.8
+27.2 / 29.3 61.9 / 64.3 45.7
+27.5 / 30.0 61.7 / 64.1 45.8
+27.8 / 30.3 60.7 / 63.4 45.6
+35.7 / 38.2 54.1 / 56.6 46.2
+Figure 7: Ablation study on each component on VG [20].
+IN: Replace the intrinsic features.
+Ex-fg: Mix up the extrinsic features of foreground triplets.
+Ex-bg: Mix up the extrinsic features of background triplets.
 Tasks.
-We evaluated models in three tasks Xu et al.
-(2017): 1) Predicate Classification (PredCls): Predicting the predicate category given all ground-truth entity bboxes and categories.
+We evaluated models in three tasks [40]: 1) Predicate Classification (PredCls): Predicting the predicate category given all ground-truth entity bboxes and categories.
 2) Scene Graph Classification (SGCls): Predicting categories of the predicate and entity given all ground-truth entity bboxes.
@@
 3) Mean: It is the average of all R@K and mR@K scores.
-Since R@K favors head predicates while mR@K favors tail predicates, the Mean can better reflect the overall performance of all predicates Li et al.
-(2022a).
-Datasets.
-We conducted all experiments on two datasets: VG and GQA.
-1) VG: We followed the widely-used strategy Xu et al.
-(2017) to split dataset (i.e., 75K/32K images for train/test, 150 object categories and 50 predicate categories).
-Besides, 5K images within the training set are sampled as val set following Tang et al.
-(2020).
-All predicate categories are divided into three groups {head, body, tail} based on the number of samples as Liu et al.
-(2019).
-2) GQA: We followed the prior work Dong et al.
-(2022) to split dataset (i.e., 70%/30% of the images for train/test, 200 object categories and 100 predicate categories).
-Similarly, 5K images within the training set are sampled as val set.
-We utilized the same group category ratio as VG to divide all predicate categories into {head, body, tail} groups.
-Implementation Details.
-Following the training protocol in prior SGG works Tang et al.
-(2020), we adopted the object detector Faster R-CNN with the ResNeXt-101-FPN Lin et al.
-(2017) backbone trained by Tang et al.
-(2020) to detect all the bounding boxes and extract their visual features.
-The parameters of the backbone were kept frozen during the training.
-This detector could achieve 28.14 mAP on the VG test set (i.e., using 0.5 IoU threshold for evaluation).
-To avoid messy hyperparamter tuning, most of the hyperparamters follow preivous works.
-More specifically, the hyperparameters $\lambda$ and $\gamma$ were set to 0.07 and 0.7 in MCFA (cf., Sec.
-3.2).
-The weights of pattern, context and semantic similarity were 1.0, 1.0 and 0.01 in intrinsic-CFA (cf., Sec.
-3.2.1).
-The hyperparameter $\alpha$ used to control the value range of adaptive mixup coefficient $\theta$ was set to 0.5 in extrinsic-CFA (cf., Sec.
-3.2.2) The threshold $\sigma$ was set to 0.5 in spatial restriction (cf., Sec.
-3.2.2).
-The $\beta$ was set to 0.1 to regulate the loss during training (cf., Sec.
-3.4).
-In this paper, SGD optimizer was used to train the model.
-The batch size was set to 12 and the initial learning rate was set to 0.01.
-After the performance on the val set reached the plateau period, the learning rate would be decayed by 10 for two times.
-All experiments were carried out with PyTorch and NVIDIA 2080Ti GPU.
+Since R@K favors head predicates while mR@K favors tail predicates, the Mean can better reflect the overall performance of all predicates [21].
+Datasets and Implementation Details. We conducted all experiments on two datasets: VG [20] and GQA [16]. More details of datasets and implementation are in the appendix.
 4.2 Comparison with State-of-the-Arts
 Setting.
-Due to the model-agnostic nature, we equipped our MCFA with three strong two-stage SGG baselines: Motifs Zellers et al.
-(2018), VCTree Tang et al.
-(2019) and Transformer Vaswani et al.
-(2017), and they are denoted as Motifs+MCFA, VCTree+MCFA, and Transformer+MCFA, respectively.
-In addition, due to the limited diversity of tail predicate components Li et al.
-(2022d), it has a high correlation with the category of subject & object.
+Due to the model-agnostic nature, we equipped our CFA with three strong two-stage SGG baselines: Motifs [44], VCTree [36] and Transformer [37], and they are denoted as Motifs+CFA, VCTree+CFA, and Transformer+CFA, respectively.
+In addition, due to the limited diversity of tail predicate components [28], it has a high correlation with the category of subject & object.
 Thus, we further equip our three models with the component prior knowledge collected from the dataset to further improve the performance of the tail predicates.
-And they are denoted as Motifs+MCFA, VCTree+MCFA, and Transformer+MCFA (More details about the priors are discussed in appendix).
+And they are denoted as Motifs+CFA, VCTree+CFA, and Transformer+CFA (More details about the priors are discussed in appendix).
 Baselines.
-We compared our methods with the state-of-the-art (SOTA) models in the VG dataset (Table 1 and Table 2) and the GQA dataset (Table 3).
-Specifically, these models can be divided into three groups: 1) Model-specific designs: Motifs, VCTree, Transformer, and BGNN Li et al.
-(2021b).
-2) Model-agnostic trade-off methods: consider the performance of all predicates comprehensively (i.e., higher Mean), e.g., PCPL Yan et al.
-(2020), DLFE Chiou et al.
-(2021), BP-LSA Guo et al.
-(2021), NICE Li et al.
-(2022a), IETrans Zhang et al.
-(2022), CFA Li et al.
-(2023a), and DPL Jeon et al.
-(2024).
-3) Model-agnostic tail-focused methods: improve tail predicates performance at the expense of excessively sacrificing the head (i.e., higher mR@K and R@50 less than 50.0% on PredCls), e.g., TDE Tang et al.
-(2020), CogTree Yu et al.
-(2021), RTPB Chen et al.
-(2022), PPDL Li et al.
-(2022c), GCL Dong et al.
-(2022), and HML Deng et al.
-(2022).
+We compared our methods with the SOTA models in the VG dataset (Table 1, Table 2) and the GQA dataset (Table 3).
+Specifically, these models can be divided into three groups: 1) Model-specific designs: Motifs, VCTree, Transformer, and BGNN [26].
+2) Model-agnostic trade-off methods: consider the performance of all predicates comprehensively (i.e., higher Mean), e.g., PCPL [41], DLFE [5], BP-LSA [11], NICE [21], and IETrans [45].
+3) Model-agnostic tail-focused methods: improve tail predicates performance at the expense of excessively sacrificing the head (i.e., higher mR@K and R@50 less than 50.0% on PredCls), e.g., TDE [35], CogTree [42], RTPB [1], PPDL [27], GCL [10], and HML [7].
 For a fair comparison, we compared with the methods in the last two groups.
-Figure 7: The examples of reasonable entities for the query triplet in intrinsic-CFA.
-The lefts are original entities of the query triplets.
-The rights are some alternate entity categories (red) and their samples (gray boxes) for each query triplet to replace.
-Figure 8: The examples of generated triplets for image-level augmentation.
 Quantitative Results on VG.
-Analysis of the trade-off methods presented in Table 1 highlights two key observations: 1) Compared to three strong baselines (i.e., Motifs, VCTree, and Transformer), MCFA significantly enhances performance on the mR@K metric across all three experimental settings.
-2) MCFA strikes the most effective trade-off between R@K and mR@K, evidenced by the highest Mean score.
-Notably, it surpasses the previous SOTA trade-off method, CFA Li et al.
-(2023a), which itself adeptly minimized performance decline on head predicates (R@K) while preserving tail performance (mR@K), in this crucial Mean metric under all settings.
-Turning to the tail-focused methods detailed in Table 2, MCFA, which incorporates further strategies dedicated to improving tail predicate performance, achieves the highest mR@K values.
-It outperforms prominent SOTA tail-focused approaches, i.e., HML Deng et al.
-(2022) and CFA Li et al.
-(2023a), on the mR@K metric across all evaluated settings.
-Additional experimental analyses are provided in the appendix.
+From the results of trade-off methods in Table 1, we can observe that: 1) Compared to the three strong baselines (i.e., Motifs, VCTree and Transformer), CFA can significantly improve model performance on mR@K metric over all three settings.
+2) CFA can achieve the best trade-off between R@K and mR@K, i.e., highest Mean, and surpass the SOTA trade-off method NICE [21] in Mean metric under all settings.
+CFA shows minimal performance degradation on the head predicates (c.f.
+R@K) while maintaining the performance of the tail (c.f.
+mR@K), demonstrating the superiority of CFA considering all predicates.
+From the results of tail-focused methods in Table 2, we can observe that: after further implementing strategies to improve tail performance, CFA can achieve the highest mR@K and exceed the SOTA tail-focused method HML [7] on mR@K metric under all settings.
+More experiment analysis is in the appendix.
 Quantitative Results on GQA.
-As demonstrated in Table 3, MCFA also markedly enhances mR@K performance over three strong baselines (i.e., Motif, VCTree, and Transformer) on the large-scale GQA dataset.
-These results certify the generalizability and effectiveness of our method.
+From the results of Table 3, we can observe that: CFA can also greatly improve the mR@K of three strong baselines (i.e., Motif, VCTree and Transformer) on the large dataset GQA, which proves the universality and effectiveness of our method.
 4.3 Ablation Studies
-Component SGCls
-IN EX-fg EX-bg mR@50 / 100 R@50 / 100 Mean
-8.7 / 9.3 39.0 / 39.7 24.2
-11.6 / 12.4 37.4 / 38.6 25.0
-15.6 / 16.5 34.9 / 36.2 25.8
-14.3 / 15.7 35.9 / 37.0 25.7
-17.0 / 18.0 35.8 / 36.9 26.9
-17.1 / 18.3 35.4 / 37.0 27.0
-17.8 / 18.8 34.4 / 35.6 26.7
-17.9 / 19.4 35.5 / 36.7 27.4
-Table 4: Ablation study on each component in feature-level augmentation Krishna et al.
-(2017).
-Motifs Zellers et al.
-(2018) is used in all ablations on VG Krishna et al.
-(2017).
-IN: Replace the intrinsic features.
-Ex-fg: Mix up the extrinsic features of foreground triplets.
-Ex-bg: Mix up the extrinsic features of background triplets.
-Similarity SGCls
+Similarity PredCls
 Pattern Context Semantic mR@50 / 100 R@50 / 100 Mean
-16.1 / 17.3 34.4 / 35.8 25.9
-17.1 / 18.3 36.1 / 37.3 27.2
-18.1 / 19.1 33.5 / 34.8 26.4
-16.9 / 18.2 35.0 / 36.3 26.6
-17.4 / 18.5 35.7 / 36.8 27.1
-16.5 / 17.6 37.0 / 38.2 27.3
-17.9 / 19.4 35.5 / 36.7 27.4
-Table 5: Ablation study on each similarity in clustering of intrinsic-MCFA.
-Feature-Level Augmentation.
-We assessed the impact of each component within our feature-level CFA using the Motifs Zellers et al.
-(2018) baseline under the SGCls setting.
-Feature-level CFA comprises three primary components: replacing intrinsic features (IN), mixing extrinsic features of foreground triplets (EX-fg), and mixing extrinsic features of background triplets (EX-bg).
-The results, reported in Table 8, lead to the following observations: 1) Employing only the IN component yields a marginal improvement in mR@K (e.g., 2.9%–3.1% gains) while slightly degrading R@K performance (a 1.1%–1.6% loss).
-This is likely attributable to the limited diversity among tail predicate triplets, meaning that feature variety remains constrained even after intrinsic feature replacement.
-2) Both EX-fg and EX-bg components substantially enhance mR@K and also achieve R@K scores relative to the baseline.
-For instance, mR@100 increases to 18.0% from the baseline’s 9.3%, and R@100 maintains 18.8% with these components.
-3) Combining all three components (IN, EX-fg, and EX-bg) achieves the optimal trade-off, i.e., the highest Mean score.
-Similarity of Clustering in Intrinsic-CFA.
-We investigated the impact of three distinct similarity measures (pattern, context, and semantic) on clustering within intrinsic-CFA, utilizing the Motifs Zellers et al.
-(2018) model as a baseline under the SGCls setting.
-Results presented in Table 5 demonstrate that relying solely on pattern similarity yields the poorest SGG performance.
-This deficiency occurs since pattern-based similarity alone disregards crucial contextual and semantic information, which is essential for guiding meaningful cluster formation.
-For instance, while both a boat and a car may share a “ridable” pattern, substituting a car for a boat would be semantically invalid in a maritime context (e.g., on the sea), as a car cannot operate there.
-Conversely, when contextual and semantic similarities are integrated into the analysis, SGG performance becomes substantially more robust, culminating in the highest Mean.
-$K_{p}$ SGCls
+33.4 / 36.2 52.0 / 54.0 43.9
+34.6 / 37.3 55.2 / 57.0 46.0
+35.3 / 38.0 51.4 / 54.1 44.7
+35.3 / 37.8 52.6 / 55.2 45.2
+35.2 / 37.9 54.0 / 56.2 45.8
+34.5 / 36.9 55.2 / 57.5 46.0
+35.7 / 38.2 54.1 / 56.6 46.2
+$K$ PredCls
 mR@50 / 100 R@50 / 100 Mean
-10 17.5 / 18.9 35.3 / 37.1 27.2
-20 17.9 / 19.4 35.5 / 36.7 27.4
-50 17.3 / 18.4 35.5 / 37.3 27.1
-Table 6: Ablation study on the number of predicate clusters $K_{p}$ of intrinsic-CFA.
-Different Cluster Number.
-We performed $K_{p}$in {10, 20, 50} to evaluate the impact of the number of predicate clusters under SGCls setting with Motifs Zellers et al.
-(2018).
-As illustrated in Li et al.
-(2023a), since when entity cluster number $K_{e}$ set to 15, there is a larger selection range of replaceable entity categories and richer feature diversity of tail predicate triplets, we use same entity cluster number setting.
-All results are reported in Table 6.
-We find that $K_{p}$ = 20 works best in mR@K.
-The reason may be that when $K_{p}$ is smaller, such as 10, the predicate clusters might be too coarse, grouping semantically distinct predicates together.
-This over-generalization can lead to less precise entity selection by potentially merging entities that play truly different roles across these distinct predicates, thereby reducing the discriminativeness for specific relations.
-When $K_{p}$ is larger (e.g., 50), the predicate clusters may become overly granular, which could prevent the model from capturing broader semantic similarities between closely related predicates.
-Consequently, this limits the effectiveness of relation-conditioned entity merging as each predicate cluster becomes too specific and offers fewer possibilities for generalization.
-$\theta$ SGCls
+15 35.7 / 38.2 54.1 / 56.6 46.2
+40 33.5 / 36.0 55.3 / 57.8 45.7
+150 33.2 / 35.7 55.7 / 57.9 45.6
+$\theta$ PredCls
 mR@50 / 100 R@50 / 100 Mean
-0.0 16.9 / 17.9 35.2 / 36.5 26.6
-0.5 17.1 / 18.6 36.0 / 37.2 27.2
-1.0 11.6 / 12.4 37.4 / 38.6 25.0
-Adap 17.9 / 19.4 35.5 / 36.7 27.4
-Table 7: Ablation study on mixup parameter $\theta$ of extrinsic CFA. Motifs Zellers et al. (2018) is used in all ablation on VG Krishna et al. (2017).
-Mixup Parameter $\theta$ in Extrinsic-CFA.
+0.0 27.1 / 29.5 62.2 / 64.3 45.8
+0.3 28.5 / 31.0 60.8 / 63.3 45.9
+0.5 35.7 / 38.2 54.1 / 56.6 46.2
+0.7 32.9 / 35.2 50.4 / 53.0 42.9
+1.0 19.3 / 21.2 64.7 / 66.6 43.0
+Table 4: Ablation studies on the different hyperparameters of each component of CFA. Motifs [44] is used in all ablation on VG [20].
+Effectiveness of Each Component.
+We evaluated the importance of each component of CFA based on Motifs [44] under the PredCls setting.
+There are three components of CFA: replace the intrinsic features (IN), mix up the extrinsic features of the foreground triplets (EX-fg), and mix up the extrinsic features of the background triplets (EX-bg).
+As reported in Table 7, we have the following observations: 1) Using only IN can slightly improve mR@K (e.g., 2.8% $\sim$ 3.4% gains) and slightly hurts R@K (0.6% $\sim$ % 0.9 loss).
+The reason is that there are not enough tail predicate triplets, resulting in the feature diversity still being limited after replacing the intrinsic features.
+2) Both Ex-fg and Ex-bg can significantly improve mR@K and keep competitive R@K compared to baseline (e.g., 17.8% vs.
+30.3% in mR@100, and 67.2% vs.
+63.4% in R@100).
+3) Combining all components allows for the best trade-off, i.e., the highest Mean.
+Similarity of Clustering in Intrinsic CFA.
+We analyzed the influence of three kinds of similarity (pattern, context, and semantic similarities) under the PredCls setting with baseline model Motifs [44].
+From Table 4(a), we can observe that using only pattern similarity shows the worst performance, since it ignores the contextual and semantic information which gives crucial guidance for selection.
+For example, both boat and car can be ridden, but we cannot replace boat with car because car can’t appear in the sea.
+Once the other two similarities are fused, the SGG performance becomes robust and the highest Mean is achieved.
+Different Clusters $K$ of Clustering in Intrinsic CFA.
+We performed $K$ $\in$ {15, 40, 150} to evaluate the impact of the number of clusters under PredCls setting with Motifs [44].
+All results are reported in Table 4(b).
+We find that KK=15 works best in mR@K.
+The reason may be that when KK=15, there is a larger selection range of replaceable entity categories and richer feature diversity of tail predicate triplets.
+Mixup Parameter $\theta$ in Extrinsic CFA.
 As mentioned in Sec.
 3.2.2, $\theta$ indicates the proportion of selected features in final augmented features.
-We investigated $\theta$in {0.0, 0.5, 1.0, adap} under the SGCls setting with Motifs Zellers et al.
-(2018) in Table 7.
+We investigated $\theta$ $\in$ {0.0, 0.3, 0.5, 0.7, 1.0} under the PredCls setting with Motifs [44] in Table 4(c).
 When $\theta$ is too small, the query triplet has too much impact on the other triplets in the image of the context triplet, and when $\theta$ is too large, feature augmentation of the query triplet is not strong enough.
-Our adaptive mixup operation alleviates this by dynamically calibrating $\theta$ based on the semantic similarity between the query and context predicates.
-Such calibration facilitates the targeted enrichment of tail predicates using relevant contexts, while concurrently preventing feature distortion from less pertinent ones, ultimately enhancing robustness.
-Effectiveness of Each Enhanced Strategy in MCFA.
-We verified the effectiveness of each enhanced strategy on the original CFA, as presented in Table 8.
-Key observations are as follows: 1) Employing relation-conditioned entity clustering effectively identifies more suitable entities, leading to an improvement in Recall while maintaining a comparable mean Recall.
-2) The adaptive mixup operation enables tail instances to better leverage contextual information from head instances, thereby significantly boosting mean Recall.
-3) The adoption of image-level augmentation can simultaneously enhance recall and mean recall, which proves the rationality of generated triplets.
-4) Combining all strategies allows MCFA to achieve the best Mean performance.
-Component SGCls
-En-IN En-EX IMG mR@50 / 100 R@50 / 100 Mean
-17.0 / 18.4 34.9 / 36.1 26.6
-16.4 / 17.6 37.1 / 38.2 27.3
-18.3 / 20.1 34.1 / 35.4 27.0
-18.1 / 18.9 35.0 / 36.2 27.1
-17.9 / 19.4 35.5 / 36.7 27.4
-Table 8: Ablation study on each enhanced strategy in MCFA on the VG dataset Krishna et al.
-(2017).
-En-IN: Utilize relation-conditioned entity clustering in intrinsic-CFA.
-En-EX: Employ the adaptive feature augmentation in extrinsic-CFA.
-IMG: Use the image-level augmentation.
-5 Qualitative Analysis
-Visualization of Reasonable Entities in Intrinsic-CFA.
-Fig.
-7 shows some alternative entities for query triplets.
-Although the visual features of these candidate entities can differ significantly from the original, they remain reasonable substitutions within the query triplet.
-For example, while zebra samples may vary considerably in color and texture from a giraffe, both are suitable because they share the common behavior of “eating plants”.
-Replacing a giraffe with a zebra in such instances introduces new visual information, thereby enriching the feature diversity associated with the eating predicate.
-Visualization of Generated Triplets for Image-level Augmentation.
-Fig.
-8 displays some generated triplets (cf., Sec.
-3.3), illustrating our method’s capability to generate diverse examples (e.g., handle-on back of-toilet) and to utilize the original query triplet’s bounding box for their supervision.
-6 Conclusion
+To better trade-off the performance on different predicates, we set $\theta$ to $0.5$.
+Visualization of Reasonable Entities in Intrinsic CFA.
+Figure 7 shows some alternative entities for query triplets.
+The visual features of the candidate entities are quite different from the original entity, but are still reasonable for the query triplet, e.g., the samples of zebra differ greatly in color and texture from the giraffe, but they can all eat plants.
+Replacing the giraffe with the zebra can provide new features to enrich the feature diversity of eating.
+5 Conclusion and Future Work
 In this paper, we revealed the drawbacks of existing re-balancing methods and discovered that the key challenge for unbiased SGG is to learn proper decision boundaries under the severe long-tailed predicate distribution.
-To address this, we proposed a model-agnostic MCFA framework to enrich the feature space of tail categories.
-It augments both intrinsic and extrinsic features of relation triplets from the feature-level and image-level.
-At the feature-level, a relation-conditioned entity clustering strategy was designed for intrinsic augmentation and an adaptive mix-up strategy for extrinsic augmentation, dynamically adjusting based on predicate similarity.
-For image-level augmentation, we devised a generative method leveraging powerful models like GLIGEN to create new visual appearances for triplets, thereby overcoming in-domain data limitations.
-Comprehensive experiments on the challenging VG and GQA datasets consistently demonstrated that MCFA’s effectiveness for unbiased SGG.
-Data Availability All experiments are conducted on publicly available datasets; see the references cited.
-Appendix
-This appendix is organized as follows:
-  • The hierarchical clustering process and results mentioned in Sec. 3.2.1 are shown in Sec. A.
-  • A more detailed discussion of prior knowledge mentioned in Sec. 4.2 is shown in Sec. B.
-  • More experimental comparisons and analyses are shown in Sec. C
-Input: Predicate category set $\mathcal{C}_{P}=\{c_{p_{i}}\mid i=1,2,\ldots,N_{p}\}$;
-Entity category set $\mathcal{C}_{E}=\{c_{e_{i}}\mid i=1,2,\ldots,N_{e}\}$;
-Predicate cluster similarity function $Sim_{P}$;
-Entity cluster similarity function $Sim_{E}$;
-Target number of predicate clusters $K_{p}$;
-Target number of entity clusters $K_{e}$ within each predicate cluster;
-Output: Set of all predicate clusters $\mathcal{S}^{p}=\{s^{p}_{i}|i=1,2,\cdots,K_{p}\}$
-Set of all entity clusters $\mathcal{S}^{e}=\{s^{e}_{i}|i=1,2,\cdots,K_{e}\}$ within each predicate cluster $s^{p}_{i}$
-/* --- Step 1: Predicate Clustering --- */
-/* Initialize each predicate category as a cluster */
-for $i=1,2,\cdots,N_{p}$ do
-       $s_{i}^{p}=\{c_{p_{i}}\}$
-/* Initialize the size of each cluster $l^{p}_{i}$ and the similarity matrix $A^{p}_{sim}$/
-for $i=1,2,\cdots,N_{p}$ do
-       $l^{p}_{i}=1$
-      for $j=1,2,\cdots,N_{p}$ do
-             $A_{sim}^{p}(i,j)=Sim_{P}(s^{p}_{i},s^{p}_{j})$
-            $A_{sim}^{p}(j,i)=A^{p}_{sim}(i,j)$
-/* Merge the two most similar predicate cluster until the number of clusters is smaller than $K_{p}$/
-while $\text{LEN}(\mathcal{S}^{p})>K_{p}$ do
-       $s^{p}_{i},s^{p}_{j}=\text{SELECT\_MAX}(A^{p}_{sim}(i,j)/(l^{p}_{i}+l^{p}_{j}))$
-      $s^{p}_{i}=\text{MERGE}(s^{p}_{i},s^{p}_{j})$
-      $\mathcal{S}^{p}=\mathcal{S}^{p}-{s^{p}_{j}}$
-      $l^{p}_{i}=l^{p}_{i}+l^{p}_{j}+1$
-      $\text{UPDATE}(A^{p}_{sim})$
-/* --- Step 2: Predicate Merging --- */
-for each predicate cluster $s^{p}_{i}\in\mathcal{S}^{p}$ do
-       for $c_{e_{i}}\in C_{E}$ do
-             if $c_{e_{i}}$ is in the same triplet category with $c_{p_{i}}\in s^{p}_{i}$ then
-                   $C^{s^{p}_{i}}_{E}\leftarrow c_{e_{i}}$
-/* --- Step 3: Entity Clustering within each Predicate Cluster --- */
-for each predicate cluster $s^{p}_{i}\in\mathcal{S}^{p}$ do
-       /* Initialize each entity category as a cluster */
-      for $i=1,2,\cdots,|C^{s^{p}_{i}}_{E}|$ do
-             $s_{i}^{e}=\{c_{e_{i}}\}$
-      /* Initialize the size of each cluster $l^{e}_{i}$ and the similarity matrix $A^{e}_{sim}$/
-      for $i=1,2,\cdots,N_{e}$ do
-             $l^{e}_{i}=1$
-            for $j=1,2,\cdots,N_{e}$ do
-                   $A_{sim}^{e}(i,j)=Sim_{E}(s^{e}_{i},s^{e}_{j})$
-                  $A_{sim}^{e}(j,i)=A^{e}_{sim}(i,j)$
-      /* Merge the two most similar entity cluster until the number of clusters is smaller than $K_{e}$/
-      while $\text{LEN}(\mathcal{S}^{e})>K_{e}$ do
-             $s^{e}_{i},s^{e}_{j}=\text{SELECT\_MAX}(A^{e}_{sim}(i,j)/(l^{e}_{i}+l^{e}_{j}))$
-            $s^{e}_{i}=\text{MERGE}(s^{e}_{i},s^{e}_{j})$
-            $\mathcal{S}^{e}=\mathcal{S}^{e}-{s^{e}_{j}}$
-            $l^{e}_{i}=l^{e}_{i}+l^{e}_{j}+1$
-            $\text{UPDATE}(A^{e}_{sim})$
-return $\mathcal{S}^{e}$ for each predicate cluster $s^{p}_{i}$;
-Algorithm 1 Relation-Conditioned Entity Clustering
-Figure 9: The results of relation clustering on VG Krishna et al. (2017).
-Figure 10: The results of entity clustering on VG Krishna et al. (2017).
-Appendix A Relation-Conditioned Entity Clustering
-As mentioned in Sec.
-3.2.1, we use relation-conditioned entity clustering to mine the replaceable entity categories.
-The specific implementation is detailed in Algorithm 1, which involves two main hierarchical clustering stages: first for predicates, and subsequently for entities, conditioned on these predicate clusters.
-Both stages utilize a similar iterative merging procedure.
-The cluster similarity functions, $Sim_{P}$ for predicates and $Sim_{E}$ for entities, are the weighted sum of pattern, context, and semantic similarities.
-In practice, we treat clusters as a collection of multiple entity categories, and the pattern and context similarity can be transferred directly from class-level to cluster-level.
-For semantic similarity, we take the mean similarity for the categories between clusters.
-After initializing the variables, the predicate and entity clustering stages share a similar logic in each iteration.
-First, we select the two most similar clusters (SELECT_MAX) based on similarity matrix ($A^{p}_{sim}$ for predicates or $A^{p}_{sim}$ for entities) and size of cluster ($l^{p}$ or $l^{e}$).
-Given that the distribution of different relation triplets in datasets is often highly imbalanced, we follow Wei et al.
-(2020) to employ the cluster size as a penalty term.
-This aims to prevent categories (whether predicate or entity) with a large proportion of samples from unduly dominating the clustering process.
-Subsequently, we merge the predicate or entity categories in cluster $s^{p}_{j}$ or $s^{e}_{j}$ into $s^{p}_{i}$ or $s^{e}_{i}$ (MERGE), remove cluster $s^{p}_{j}$ / $s^{e}_{j}$, and update the corresponding cluster size.
-Since the merged cluster has changed, we recalculate the similarity of other clusters to it and then update $A^{p}_{sim}$ / $A^{e}_{sim}$ (UPDATE).
-The above steps are repeated until the number of clusters is equal to $K^{p}$ for predicates or $K^{e}$ for entities.
-The predicate clustering result and entity clustering results are shown in Fig.
-9 and Fig.
-10, respectively.
-Obviously, some predicate and entity categories that share common pattern are grouped into the same cluster, e.g., lying on and laying on (predicates), and dog and cat (entities, grouped within relevant predicate cluster contexts).
-Figure 11: Performance(%) comparison between Motifs Zellers et al.
-(2018) and Motifs+CFA over all predicates on test set of VG Krishna et al.
-(2017).
-The orange area denotes the predicate distribution of training set.
-Appendix B Component Prior Knowledge
-As discussed in prior SGG works Li et al.
-(2022d), the statistic prior of the predicate distribution under a given condition can improve the mR@K performance of the unbiased SGG.
-In the inference phase, we calculate statistic component prior $b_{s,o,r}$, and add it to the predicted logits to predict predicate category:
-${{{b}}_{s,o,r}}=-\log\frac{{{count_{s,o,r}}}}{{\sum\nolimits_{i=1}^{H}{{count_{s,o,i}}}}},$ (19)
-where $H$ is the number of predicate categories.
-As for the prior “Triplet”, “Subject”, and “Object”, $count_{s,o,r}$ is the amount of the triplets whose predicate category is $r$ in the training set given subject-object pair, subject, and object respectively.
-Appendix C Extra Comparisons and Analyses
-Strategy PredCls
-mR@50 / 100 R@50 / 100 Mean
-Motifs Zellers et al. (2018) 16.5 / 17.8 65.6 / 67.2 41.8
-+Reweight Tang et al. (2020) 30.8 ↑14.3 / 34.5 ↑16.7 36.1 / 40.4 35.5
-+Resample Tang et al. (2020) 18.5 ↑2.0 / 20.0 ↑2.2 64.6 / 66.7 42.5
-+MCFA 35.8 ${}_{{\color[rgb]{0.1725,0.7059,0.3921}\definecolor[named]{pgfstrokecolor}{rgb}{0.1725,0.7059,0.3921}\uparrow\textbf{18.7}}}$ / 38.3 ${}_{{\color[rgb]{0.1725,0.7059,0.3921}\definecolor[named]{pgfstrokecolor}{rgb}{0.1725,0.7059,0.3921}\uparrow\textbf{20.5}}}$ 54.3 / 56.8 46.3
-Table 9: Performance (%) of re-balancing strategies and MCFA on VG Krishna et al. (2017).
-Figure 12: The results of scene graphs generated by Motifs (blue) and Motifs-MCFA (green) on VG Krishna et al.
-(2017).
-Green predicates are correct (i.e., match GT), brown predicates are acceptable (i.e., does not match GT but still reasonable), and purple predicates are more informative (i.e., does not match GT and more reasonable).
-Figure 13: The t-SNE visualization of randomly sampled instances of two predicates on the feature space before and after MCFA.
-C.1 Comparison with Re-balancing Methods.
-To demonstrate the superiority of MCFA compared with the prevalent re-balancing methods (i.e., reweight and resample) Tang et al.
-(2020), we conducted the three strategies on the baseline Motifs Zellers et al.
-(2018).
-The results under the PredCls setting are reported in Table 9.
-From the results, we can observe: 1) The Motifs baseline can achieve the best R@K.
-However, the high R@K is mainly due to the frequency bias of the dataset Tang et al.
-(2020), and they suffer severe drops in tail predicates.
-2) The reweight method achieves better performance at mR@K, but it also sacrifices the performance of head predicates excessively, resulting in a low Mean.
-3) The resample method keeps R@K high, but the improvement of mR@K is slight.
-The reason is that those decision boundaries may still be biased toward the head.
-4) MCFA achieves the highest mR@K and maintains high R@K, i.e., it achieves the best trade-off over different predicate categories (highest performance on Mean).
-C.2 Comparison over All Predicates
-To further demonstrate the performance of each predicate, we displayed the predicate distribution of the training set in the VG dataset Krishna et al.
-(2017) and the performance of Motifs Zellers et al.
-(2018) and Motifs+MCFA on R@100 for each predicate category in Fig.
-11.
-Obviously, our approach slightly compromises the performance of the head predicates (e.g., wearing), but greatly improves the tail predicates (e.g., flying in).
-This proves the superiority of our method in considering the performance of all predicates.
-C.3 Qualitative Results.
-Figure 12 shows some qualitative results generated by Motifs Zellers et al.
-(2018) and Motifs+MCFA under PredCls setting.
-We can observe that MCFA can not only predict more accurate predicates (e.g., on vsparked on and on vs.walking on), but also more fine-grained and informative predicates (e.g., holding vsusing).
-C.4 Quality of Compositional Features.
-We visualized the t-SNE distributions of features (from the previous layer of classification) in Fig.
-13, demonstrating that our MCFA enhances the distinctiveness of tail predicates in feature space, i.e., it aids in finding suitable decision boundaries and does not harm to training, resulting in improved performance.
+Thus, we proposed a model-agnostic CFA framework that can enrich the feature space of tail categories by augmenting both intrinsic and extrinsic features of the relation triplets.
+Comprehensive experiments on the challenging VG and GQA datasets showed that CFA significantly improves the performance of unbiased SGG.
+In the future, we would like to extend CFA to compose new triplets features by fusing features from open-vocabulary categories or images from other domains.
+Acknowledgement.
+This work was supported by the National Key Research & Development Project of China (2021ZD0110700), the National Natural Science Foundation of China (U19B2043, 61976185), and the Fundamental Research Funds for the Central Universities(226-2023-00048).
+Long Chen was supported by HKUST Special Support for Young Faculty under Grant F0927.